### v3

# Benchmarking Vision Foundation Models for Input Monitoring in Autonomous Driving

Abstract
Deep neural networks (DNNs) remain challenged by distribution shifts in complex open-world domains like automated driving (AD): Robustness against yet unknown novel objects (semantic shift) or styles like lighting conditions (covariate shift) cannot be guaranteed. Hence, reliable operation-time monitors for identification of out-of-training-data-distribution (OOD) scenarios are imperative. Current approaches for OOD classification are untested for complex domains like AD, are limited in the kinds of shifts they detect, or even require supervision with OOD samples. To prepare for unanticipated shifts, we instead establish a framework around a principled, unsupervised and model-agnostic method that unifies detection of semantic and covariate shifts: Find a full model of the training data’s feature distribution, to then use its density at new points as in-distribution (ID) score. To implement this, we propose to combine Vision Foundation Models (VFMs) as feature extractors with density modeling techniques. Through a comprehensive benchmark of 4 VFMs with different backbone architectures and 5 density-modeling techniques against established baselines, we provide the first systematic evaluation of OOD classification capabilities of VFMs across diverse conditions. A comparison with state-of-the-art binary OOD classification methods reveals that VFM embeddings with density estimation outperform existing approaches in identifying OOD inputs. Additionally, we show that our method detects high-risk inputs likely to cause errors in downstream tasks, thereby improving overall performance. Overall, VFMs, when coupled with robust density modeling techniques, are promising to realize model-agnostic, unsupervised, reliable safety monitors in complex vision tasks.
1 Introduction
Figure 1: A monitoring system for autonomous driving that uses a pre-trained Vision Foundation Model’s image encoder to detect distribution shifts in input data. The system identifies whether camera inputs (red dots) fall within the model’s training distribution (brown dots) or represent covariate shifts (top) or semantic shifts (bottom), helping assess operational safety and potential failure risks.
In recent decades, deep neural networks (DNNs) have become widely used for environmental perception due to their superior performance compared to traditional computer vision methods [13, 15, 90, 58, 73, 48, 45]. However, in complex open-world domains such as autonomous driving (AD) and robot navigation this performance is inevitably challenged by the “long tail” of the input distribution, i.e., unforeseeable out-of-distribution (OOD) scenarios for which DNN generalization cannot be guaranteed. AD spans multiple operational levels, from driver assistance (L1-L2) to fully autonomous operation (L4-L5), with conditional automation (L3) bridging the transition [67]. Across these levels, camera-based perception systems support critical functions, from lane-keeping and traffic sign recognition to full scene understanding [9]. As these systems operate in increasingly diverse environments, a fundamental question emerges: How can we ensure that perception models function reliably in complex, open-world scenarios, particularly when encountering inputs that deviate from their training distribution? To mitigate these risks and ensure safe operation, reliable monitors that can detect and respond to such deviations have become mandatory [17, 34]. Aside from model uncertainty and behavior monitoring [62, 25], an important approach for OOD detection is input monitoring (Fig. 1), which identifies when input data fed to the system deviates from the training data distribution [62].
Existing binary OOD classification methods suffer from critical limitations for AD applications: they often require direct access to model internals, are specialized to specific architectures, or need exposure to anticipated shift types during training [32, 46, 51, 89, 76]. These constraints compromise reliability when encountering truly unforeseen distribution shifts in open-world environments – a critical limitation for autonomous driving systems that must handle diverse operational domains. Furthermore, accessing internal model layers becomes impractical in AD systems. A more principled approach would model the complete training data distribution in feature space, using density values or likelihoods as in-distribution (ID) scores for new samples. The challenge for complex visual domains lies in extracting sufficiently powerful features—standard DNNs and autoencoders (AE) trained on limited datasets typically fail to capture the necessary representational richness.
This is precisely where VFMs offer a compelling solution. Trained on diverse, large-scale datasets, VFMs encode comprehensive semantic contexts and subtle visual distinctions within their latent representations [10, 56], making them particularly well-suited for distribution modeling. Their appeal for input monitoring in AD stems from demonstrated strong generalization capabilities across diverse visual tasks and novel scenarios [33, 54], potentially enabling robust detection of unforeseen distribution shifts without requiring model-specific adaptations or prior exposure to OOD examples.
In this study we investigate: How suitable are VFM image encodings for modeling a data distribution for OOD scenario detection in complex vision domains like AD? In particular: How do they compare to standard DNN encodings and SoTA baseline binary-OOD detection methods? Is this OOD of use for improving overall system reliability in downstream applications? Also, while most existing works focuses on either semantic OOD detection on object level [11, 5, 12] or on covariate shifts detection [6, 29], we are interested in monitoring capabilities for both semantic, covariate, and combined shift detection on scene level. Our contributions are:
  1. (i)
    We propose a modular pipeline for unsupervised, model-agnostic detection of distribution shifts in input images (see Fig. 1), in particular both semantic, covariate, and joint shifts. At its core we use rich VFM feature encodings for distribution density modeling.
  2. (ii)
    We for the first time benchmark VFMs for OOD on the challenging open-world domain of AD. Our extensive tests compare 4 VFMs on 12 both synthetic and real datasets, and 7 OOD methods as baselines. Our results show the clear superiority of VFM encodings in unified distribution shift detection.
  3. (iii)
    We demonstrate practical usability on a downstream AD perception application, showcasing that, despite being model-agnostic, VFMs successfully identify OOD samples with higher risk of errors.
  4. (iv)
    We establish a comprehensive benchmarking framework for evaluating binary OOD classification methods in AD by providing diverse experimental scenarios combining different types of distribution shifts, datasets, and evaluation protocols.
2 Related Work
2.1 Semantic Shift in Autonomous Driving
In machine learning, training and testing sets are typically drawn from the same distribution, which poses the closed-world assumption [62] that the set of considered semantic classes will not change [59, 30]. To identify semantic shifts (novel classes), techniques like Monte Carlo dropout estimate neural network uncertainty [21], finding applications in semantic segmentation [38] and object detection [27, 28]. Recent advances integrate semantic segmentation with anomaly detection [5, 12] and employ various reconstruction approaches [79, 55, 82] for object-level OOD detection. Alternative feature-based approaches include SAFE [81], which teaches detectors to recognize sensitivity-aware features through input perturbation, and FOOD-ID [41], which employs likelihood-based feature space mechanisms. However, existing solutions typically rely on model-specific architectures, OOD training examples, computationally expensive generative models, or are limited to specific anticipatable shift types. In contrast, pre-trained VFMs enable unsupervised, model-agnostic detection of distribution shifts without these limitations.
2.2 Covariate Shift in Autonomous Driving
The distribution of real-world scenarios often diverges from training data, a phenomenon known as covariate shift [70], distributional shift [29], or domain shift [49]. In autonomous driving, this occurs due to changes in lighting conditions and weather patterns. Comparative studies [29] found likelihood regret [85] and self-supervised approaches [71] outperform variational autoencoders [74] for covariate shift detection. Other methods include Earth Mover’s Distance for linking domain mismatch with performance degradation [49], feature map error analysis [6], and GAN-based validation frameworks [88].
Similar to semantic shift detection, these approaches rely on model-specific architectures, require supervision with specific shift examples, or depend on computationally expensive generative models. Our work addresses these fundamental limitations by leveraging VFMs’ rich representational power for model-agnostic distribution shift detection that unifies both covariate and semantic shift identification.
2.3 Vision Foundation Models
FMs are large-scale machine learning models trained on diverse data, allowing them to adapt to varied tasks beyond their original objective [83]. The field includes models like CLIP [61], trained contrastively on image-text pairs; DINO [10], using self-supervised learning with self-distillation; DINOv2 [56], enhancing DINO with masked image modeling; Grounding-DINO [45], extending DETR-based detection [86] with text guidance; InternImage [80], utilizing deformable convolutions for adaptive receptive fields; Florence-2 [84], employing hierarchical visual tokenization; and LLaVA [44], connecting vision encoders to large language models through visual instruction tuning. This study focuses on the first four VFMs, selected for their architectural diversity (spanning Convolutional Networks, Vision Transformers, and Swin Transformers) and absence of test data leakage (see Supplementary). For a comprehensive overview of FMs, we refer to [7, 2], while [91] provides a detailed review of VFMs specifically in AD.
Note that, while CLIP [61] and GroundingDINO [45] may also accept language inputs, we here for now only consider their image encodings.
3 Shift Detection Methodology & Evaluation
3.1 Shift Detection Methodology: Overview
We propose a method to detect any distribution shift in image data using a DNN $F\colon x\mapsto f$ for feature extraction (a VFM, standard DNN, or AE). Concretely, we aim to classify whether a new image $x^{*}$ is an outlier with respect to the distribution of a given (monitor training) dataset $\mathcal{X}_{\mathrm{monitor}}$ (the ID data). In our case we choose $\mathcal{X}_{\mathrm{monitor}}=\mathcal{X}_{\mathrm{perception}}$ (e.g., Cityscapes train set [16]), the dataset the perception system is trained on. During monitor training, the monitor is then parametrized on the (fixed) features $\mathcal{F}_{\mathrm{monitor}}\coloneqq F(\mathcal{X}_{\mathrm{monitor}})$ to capture the distribution $\mathcal{F}_{\mathrm{monitor}}$. From the distribution model we then derive a function $c$ that provides an ID score $c(x^{*})=\mathrm{score}(f^{*})$ for any new input $x^{*}$, based on its similarity to or density in the monitor data feature distribution. Using few ID and OOD samples, a decision threshold $t$ is determined. For monitor inference on new samples $x*$, $c(x^{*})>t$ is calculated to decide the OOD vs. ID binary classification. For testing, we combine an ID dataset (e.g., Cityscapes validation set [16]) with an OOD dataset (e.g., Foggy Cityscapes [69]), both containing samples not seen by the monitor previously: $\mathcal{X}_{\mathrm{test}}=\mathcal{X}_{\mathrm{ID}}\cup\mathcal{X}_{\mathrm{OOD}}$. The OOD samples differ from $\mathcal{X}_{\mathrm{monitor}}$ through either semantic shifts (e.g., novel object classes) or covariate shifts (e.g., varying lighting conditions or adverse weather).
We consider the following choices for the monitor distribution modeling further detailed in the following in order of increasing complexity: 1) instance-based: averaged pairwise similarities to the training examples and its computationally efficient derivation using similarity to the mean feature vector, 2) a one-class support vector machine, 3) a Gaussian mixture model and 4) a normalizing flow.
Average Pairwise Similarity (APS) and Mean Feature Similarity (MFS):
In this method, we compute the similarity between a test image and training images using cosine similarity. The test set includes both $\mathcal{X}_{\mathrm{ID}}$ and $\mathcal{X}_{\mathrm{OOD}}$ images. Cosine similarity is chosen for several reasons: (1) its effectiveness in measuring the similarity between high-dimensional vectors, particularly when the magnitude of the vectors is less important than their direction; (2) its invariance to scaling, which is beneficial when dealing with normalized feature vectors; and (3) its natural alignment with the latent space of many pre-trained foundational models used for feature extraction. Notably, models like CLIP and DINO are often trained using contrastive losses based on cosine similarity, making this metric particularly suitable for comparing features in their embedding spaces. For a new test image $x^{*}$ with feature vector $f^{*}$, we calculate its average cosine similarity to all training samples.
$c_{\text{APS}}(x^{*})=\frac{1}{n}\sum_{i=1}^{n}\cos(f^{*},f_{i})=\frac{1}{n}\sum_{i=1}^{n}\frac{f^{*}\cdot f_{i}}{||f^{*}||\,||f_{i}||}$ (1)
where $\cos(f^{*},f_{i})$ denotes the cosine similarity between vectors $f^{*}$ and $f_{i}$, $\cdot$ denotes the dot product, and |||||$\cdot$||| | ⋅ | | is the L2 norm. The resulting similarity score $c(x^{*})$ ranges from -1 to 1, with higher values indicating greater similarity.
For real-time applications, we also implement MFS, a computationally efficient variant that first computes a mean feature vector $\bar{f}=\frac{1}{n}\sum_{i=1}^{n}f_{i}$ from training samples, then calculates similarity to this prototype:
$c_{\text{MFS}}(x^{*})=\cos(f^{*},\bar{f})=\frac{f^{*}\cdot\bar{f}}{||f^{*}||\,||\bar{f}||}$ (2)
This reduces inference complexity from $O(n)$ to $O(1)$, making it suitable for monitoring systems with constrained computational resources.
One-Class Support Vector Machine (OC-SVM):
OC-SVM [8, 1] is an unsupervised machine learning algorithm designed for novelty detection. It is particularly effective in scenarios where there is an abundance of data from one class (the ”normal” class) and the objective is to identify instances that deviate from this norm. In our context, we employ OC-SVM to model the distribution of features extracted from $\mathcal{X}_{\mathrm{monitor}}$ images and subsequently detect samples from $\mathcal{X}_{\mathrm{ID}}$ and $\mathcal{X}_{\mathrm{OOD}}$ sets. For a set of $\mathcal{X}_{\mathrm{monitor}}$ features $\mathcal{F}_{\mathrm{monitor}}=\{f_{1},f_{2},\ldots,f_{n}\}$ extracted from $\mathcal{X}_{\mathrm{monitor}}$ set, the OC-SVM aims to find a hyperplane that separates the data from the origin with maximum margin in the feature space. This is achieved by solving the following optimization problem:
$\min_{w,\xi,\rho}\frac{1}{2}||w||^{2}+\frac{1}{\nu n}\sum_{i=1}^{n}\xi_{i}-\rho$ (3)
subject to: $(w\cdot\phi(f_{i}))\geq\rho-\xi_{i},\quad\xi_{i}\geq 0$ (4)
where $w$ is the normal vector to the hyperplane, $\phi$ is a feature map that implicitly maps the input features to a higher-dimensional space, $\xi_{i}$ are slack variables, $\rho$ is the offset of the hyperplane from the origin, and $\nu\in(0,1]$ is a parameter that controls the trade-off between maximizing the distance of the hyperplane from the origin and the fraction of training examples that are allowed to fall on the wrong side of the hyperplane. For a new test image $x^{*}$ with feature vector $f^{*}$, the decision function is:
$\SwapAboveDisplaySkip c_{\text{OC-SVM}}(x^{*})=(w\cdot\phi(f^{*}))-\rho$ (5)
We utilize these decision function values as anomaly scores, with lower scores indicating a higher likelihood of the sample being OOD.
Gaussian Mixture Model (GMM):
We model the features $\mathcal{F}_{\mathrm{train}}$ using a Gaussian mixture model (GMM) [52] of $K$ Gaussians with the formulation
$\SwapAboveDisplaySkip p(f_{i}|\Theta)=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(f_{i}|\mu_{k},\Sigma_{k}),$ (6)
Therein, $\Theta=\{\pi_{1},\ldots,\pi_{K},\mu_{1},\ldots,\mu_{K},\Sigma_{1},\ldots,\Sigma_{K}\}$ represents the GMM parameters, where $\pi_{k}$ is the mixing coefficient (with $\sum_{k=1}^{K}\pi_{k}=1$ ensuring a valid probability distribution), $\mu_{k}$ the mean vector, and $\Sigma_{k}$ the covariance matrix of the $k$-th Gaussian component.
The parameters $\Theta$ of the GMM are derived from the feature set $\mathcal{F}_{\mathrm{train}}$ using the Expectation-Maximization (EM) algorithm. To optimally balance complexity and fitting accuracy, the model’s number of components $K$ is determined through the Akaike Information Criterion (AIC). This criterion is applied by testing a series of $K$ values on $\mathcal{F}_{\mathrm{monitor}}$. The choice of $K$ is discussed in the Appendix.
We further assume to have an ID set $\mathcal{X}_{\mathrm{ID}}$ with corresponding features $\mathcal{F}_{\mathrm{ID}}$ and analogously OOD set (containing either semantic or covariate shift) $\mathcal{X}_{\mathrm{OOD}}$ with corresponding features $\mathcal{F}_{\mathrm{OOD}}$. Let $\mathcal{X}=\mathcal{X}_{\mathrm{ID}}\cup\mathcal{X}_{\mathrm{OOD}}$ and $\mathcal{F}=\mathcal{F}_{\mathrm{ID}}\cup\mathcal{F}_{\mathrm{OOD}}$. For a new test image $x^{*}\in\mathcal{X}$, we extract and normalize its feature vector $f^{*}$ using the given foundational model currently under inspection. To compute confidence scores, the log-likelihood of $f^{*}$ under the learned GMM is computed as
$\SwapAboveDisplaySkip c_{\text{GMM}}(x^{*})=\log p(f^{*}|\Theta)=\log\left(\sum_{k=1}^{K}\pi_{k}\mathcal{N}(f^{*}|\mu_{k},\Sigma_{k})\right)$ (7)
where $\Theta=\{\pi_{k},\mu_{k},\Sigma_{k}\}_{k=1}^{K}$ is the set of parameters for the GMM, learned from $\mathcal{F}_{\mathrm{train}}$.
Normalizing Flows (NF):
Normalizing flows [57, 19] provide a deep learning framework for high-dimensional density estimation through invertible mappings. Let $P_{F}$ denote the probability measure of VFM features $f_{1},\ldots,f_{n}$ from $\mathcal{F}_{\mathrm{monitor}}$, and $P_{\mathcal{N}}$ denote the measure of the multivariate standard normal distribution $\mathcal{N}(\cdot|0,I)$. A normalizing flow learns an invertible map $g_{\theta}$ such that:
$\SwapAboveDisplaySkip P_{F}\circ g_{\theta}^{-1}\approx P_{\mathcal{N}}$ (8)
where $g_{\theta}^{-1}$ denotes the pre-image. We want to use the density corresponding to $P_{F}$ at a feature vector $f$ to derive an ID-score for $f$. Given $g_{\theta}$ as above, the density can be comfortably estimated as $\mathcal{N}(g_{\theta}(f)|0,I)\cdot|\det(J_{f}g_{\theta}(f))|$, where $J_{f}g_{\theta}$ is the Jacobian of $g_{\theta}$ with respect to $f\in\mathbb{R}^{d}$.
We implement $g_{\theta}$ using a Real-valued Non-Volume Preserving (Real-NVP) [19] architecture with affine coupling layers. Each layer employs a binary mask to split input dimensions into $x_{1}$ and $x_{2}$, producing outputs:
$y_{1}$ $=x_{1}$ $y_{2}$ $=x_{2}\odot\exp(s(x_{1}))+t(x_{1})$ (9)
Here, $s(\cdot)$ and $t(\cdot)$ are neural networks computing scale and translation respectively, and $\odot$ is the element-wise product. This architecture’s Jacobian determinant simplifies to $\log|\det(J)|=\sum_{i}s(x_{1})_{i}$, enabling efficient likelihood computation. For OOD detection, we normalize the negative log-likelihood of test samples to [0,1], where higher values indicate potential OOD samples.
4 Experiments
This section presents a comprehensive empirical evaluation, focusing on systematic experiments to assess the ability of VFMs to detect different distributional changes in the challenging AD domain. We begin by reviewing the datasets used in our experiments and provide implementation details for the methods benchmarked in this study. Our evaluation examines the detection of semantic shifts, covariate shifts, and their combination. The proposed unsupervised shift detection techniques, based on modeling the distribution of $\mathcal{F}_{\mathrm{monitor}}$, are systematically compared to state-of-the-art OOD detection methods. Additionally, we evaluate our technique using identical model architectures trained on different datasets to establish comparative baselines. This broad experimental evaluation is complemented with a downstream application: we employ our shift detection techniques to filter out OOD inputs to semantic segmentation DNNs, and show that increasing filtering strengths notably improves segmentation performance. This suggests that our model-agnostic method specifically captures error-prone OOD samples despite being model-agnostic.
Metrics:
For each test image $x^{*}\in\mathcal{X}=\mathcal{X}_{\text{ID}}\cup\mathcal{X}_{\text{OOD}}$, we compute a confidence score $c(x^{*})$ using one of the models outlined above. Higher values of $c(x^{*})$ indicate a greater likelihood that $x^{*}$ belongs to the ID data represented by the set $\mathcal{X}_{\mathrm{ID}}$. To evaluate the model’s performance in distinguishing between ID and OOD samples, we consider the following binary classification metrics:
Area Under the Receiver Operating Characteristic Curve, AUROC: The ROC curve plots true against false positive rate for varying decision thresholds. An AUROC value of 1 indicates perfect classification, while 0.5 corresponds to random guessing.
Area Under the Precision-Recall Curve, AUPR: The AUPR is particularly useful for performance evaluation on class-imbalanced test datasets.
False Positive Rate (FPR) at 95% True Positive Rate (TPR), FPR95: This metric measures the FPR for the threshold at which the TPR is 95%. Specifically, we determine the threshold $t$ that achieves a 95% TPR for ID samples and compute the corresponding FPR for OOD samples at $t$. Lower FPR95 values mean fewer OOD samples are misclassified as ID, i.e., better OOD detection performance.
These metrics are measured on $\mathcal{X}_{\text{ID}}\cup\mathcal{X}_{\text{OOD}}$.
Datasets:
Our experimental framework employs a diverse array of datasets to evaluate the robustness of our proposed method under various shifts. We carefully selected and curated datasets to simulate semantic shifts, covariate shifts, and their combination, providing a comprehensive testbed for our analysis.
To investigate semantic shifts, we use the Cityscapes train set [16] as our primary $\mathcal{X}_{\mathrm{monitor}}$, with its validation set serving as $\mathcal{X}_{\mathrm{ID}}$. We introduced OOD elements using two complementary datasets: the Lost and Found dataset [59], which captures real-world road hazards in environmental conditions matching Cityscapes, and the bravo-synobjs dataset [50], which provides synthetic anomalies through generative inpainting of Cityscapes images. This combination enables evaluation of VFMs on both naturally occurring and synthetically generated semantic shifts while maintaining consistent environmental conditions with $\mathcal{X}_{\mathrm{monitor}}$.
For covariate shifts, we employed datasets that maintain semantic consistency with $\mathcal{X}_{\mathrm{monitor}}$ while introducing varying visual characteristics. Using the Cityscapes training set [16] as our monitoring dataset $\mathcal{X}_{\mathrm{monitor}}$, validation set of Cityscapes as $\mathcal{X}_{\mathrm{ID}}$, we evaluate on several OOD datasets $\mathcal{X}_{\mathrm{OOD}}$: Foggy Cityscapes [68] with graduated synthetic fog, bravo-synrain and bravo-synflare [50] with synthetic rain and lens flare effects. For ACDC, we utilize its clear-weather training images as part of $\mathcal{X}_{\mathrm{monitor}}$, clear-weather validation images as part of $\mathcal{X}_{\mathrm{ID}}$ and its adverse-condition validation images (i.e. fog, night, rain, snow) as $\mathcal{X}_{\mathrm{OOD}}$. This setup enables systematic evaluation of VFMs under both synthetic and real-world covariate shifts.
To explore the interplay between semantic and covariate shifts, we incorporated the SegmentMeIfYouCan (SMIYC) dataset [11] and the Indian Driving Dataset (IDD) [77]. These datasets introduce both novel semantic elements and environmental variations, allowing us to assess our method under complex, real-world distributional changes. See Appendix for examples and visualizations.
Pre-trained VFMs and Implementation Details:
We employ a diverse selection of pre-trained VFMs, including CLIP [61], DINO [10], DINOv2 [56], and Grounding DINO [45], each representing a distinct paradigm in visual representation learning. CLIP [61] is evaluated across multiple architectures. Specifically, we utilize ResNet (RN) variants with varying depths (RN50, RN101) and width multipliers (RN50x4, RN50x16, RN50x64), alongside Vision Transformer (ViT) configurations with different patch sizes and model capacities (ViT-B/32, ViT-B/16, ViT-L/14) [^1] CLIP: https://github.com/openai/CLIP. DINO is examined through a range of backbones, including ViT variants with diverse patch sizes (Dino-ViT-S/16, Dino-ViT-S/8, Dino-ViT-B/16, Dino-ViT-B/8) and a ResNet variant (Dino-RN50) [^2] DINO: https://github.com/facebookresearch/dino. DINOv2 [56] is represented by Vision Transformer variants of increasing model capacity (Dinov2-ViT-S/14, Dinov2-ViT-B/14, Dinov2-ViT-L/14, Dinov2-ViT-G/14) [^3] DINOv2: https://github.com/facebookresearch/dinov2. Grounding DINO [45] employs a Swin Transformer [48] as its image encoder, building upon the DETR-based detection architecture [86]. In our experiments, we include Grounding DINO to evaluate the effectiveness of the Swin architecture [48] in comparison to ViT-based models, providing a broader perspective on the performance of different backbone architectures [^4] Grounding DINO: https://github.com/IDEA-Research/GroundingDINO.
We also compare the performance of the VFMs against their ImageNet-trained counterparts [18]. This baseline includes ResNet variants (RN50, RN101, RN152) [26], ViT architectures (ViT-B/16, ViT-L/16) [20], and Swin Transformer backbones (SwinB, SwinT) [^5] ImageNet-trained ResNets: https://github.com/pytorch/pytorch; ViT and Swin Transformers: https://github.com/huggingface/pytorch-image-models [48]. Furthermore, we trained autoencoder (AE) on the primary $\mathcal{X}_{\mathrm{monitor}}$ dataset to serve as feature extractors, offering an additional point of comparison with the VFMs. Architectural specifications, comparisons with ImageNet-trained counterparts, and implementation considerations are presented in the Appendix.
4.1 Semantic Shift
Tab. 1 shows the two top-performing backbones of each model for semantic shift detection (find the complete comparison in the Appendix). The results reveal the varying sensitivities of different VFMs and their backbones to input images containing rare semantic classes absent from the training data. Several key observations are discussed in the following.
Model Backbone Latent Space Dim. Lost and Found
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC↑ AUPR↑ FPR95↓ AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50x64 1024 0.94 0.93 0.24 0.93 0.93 0.27 0.94 0.93 0.24 0.94 0.93 0.23 0.88 0.84 0.37
ViT-B/16 512 0.88 0.86 0.38 0.86 0.85 0.41 0.89 0.88 0.43 0.91 0.89 0.33 0.91 0.88 0.34
DINO ViT-B/16 768 0.79 0.75 0.52 0.72 0.64 0.60 0.79 0.76 0.59 0.84 0.82 0.55 0.75 0.71 0.76
ViT-B/8 768 0.81 0.76 0.49 0.73 0.64 0.59 0.77 0.75 0.56 0.87 0.86 0.45 0.82 0.81 0.59
DINOv2 ViT-L/14 1024 0.89 0.87 0.43 0.91 0.88 0.37 0.84 0.82 0.68 0.82 0.77 0.52 0.83 0.78 0.54
ViT-G/14 1536 0.90 0.87 0.30 0.89 0.86 0.37 0.85 0.82 0.63 0.86 0.78 0.39 0.83 0.75 0.46
Grounding DINO SwinB 768 0.88 0.84 0.47 0.82 0.76 0.52 0.84 0.80 0.54 0.91 0.89 0.32 0.95 0.94 0.16
SwinT 768 0.85 0.80 0.53 0.79 0.73 0.60 0.83 0.79 0.59 0.92 0.90 0.28 0.96 0.95 0.16
Bravo-Synobj
CLIP RN50x64 1024 0.93 0.94 0.54 0.93 0.94 0.54 0.93 0.95 0.54 0.94 0.95 0.53 0.94 0.96 0.43
ViT-B/16 512 0.91 0.93 0.50 0.91 0.93 0.50 0.91 0.93 0.54 0.92 0.94 0.48 0.93 0.95 0.49
DINO ViT-B/16 768 0.81 0.82 0.69 0.81 0.82 0.69 0.83 0.85 0.75 0.86 0.88 0.60 0.84 0.87 0.80
ViT-B/8 768 0.82 0.84 0.64 0.82 0.84 0.64 0.83 0.86 0.70 0.86 0.89 0.59 0.88 0.91 0.65
DINOv2 ViT-L/14 1024 0.78 0.79 0.77 0.78 0.79 0.77 0.78 0.81 0.80 0.82 0.85 0.74 0.86 0.88 0.63
ViT-G/14 1536 0.79 0.80 0.73 0.79 0.80 0.73 0.81 0.83 0.69 0.82 0.85 0.71 0.86 0.89 0.69
Grounding DINO SwinB 768 0.60 0.56 0.83 0.61 0.56 0.82 0.58 0.54 0.85 0.76 0.68 0.58 0.99 0.99 0.02
SwinT 768 0.57 0.54 0.89 0.58 0.54 0.90 0.57 0.54 0.91 0.64 0.58 0.78 0.91 0.84 0.20
Table 1: Comparison of VFMs with various backbones on the Lost and Found [59] and Bravo-Synobj[50] datasets. focusing on semantic shift sensitivity. Results show AUROC↑. AUPR↑ (higher is better) and FPR95↓ (lower is better) metrics. Best values per metric in bold.
Model Backbone Latent Space Dim.
Foggy
Cityscapes
(i=0.05)
Foggy
Cityscapes
(i = 0.2)
Bravo-
Synflare
Bravo-
Synrain
ACDC Fog ACDC Night ACDC Rain ACDC Snow
FPR95 FPR95 FPR95 FPR95 FPR95 FPR95 FPR95 FPR95
GMM NF GMM NF GMM NF GMM NF GMM NF GMM NF GMM NF GMM NF
CLIP RN50x64 1024 0.76 0.67 0.26 0.21 0.62 0.47 0.01 0.04 0.06 0.04 0.13 0.23 0.26 0.22 0.05 0.03
ViT-B/16 512 0.57 0.48 0.18 0.10 0.43 0.35 0.20 0.12 0.07 0.07 0.34 0.16 0.27 0.25 0.05 0.05
DINO ViT-B/16 768 0.72 0.74 0.31 0.34 0.37 0.57 0.17 0.03 0.43 0.50 0.33 0.23 0.53 0.68 0.43 0.48
ViT-B/8 768 0.72 0.52 0.35 0.09 0.51 0.33 0.10 0 0.35 0.35 0.48 0.27 0.52 0.58 0.38 0.34
DINOv2 ViT-L/14 1024 0.92 0.88 0.84 0.70 0.88 0.80 0.56 0.37 0.69 0.59 0.48 0.39 0.83 0.67 0.62 0.38
ViT-G/14 1536 0.89 0.83 0.73 0.54 0.88 0.80 0.61 0.35 0.63 0.46 0.31 0.31 0.71 0.58 0.49 0.28
Grounding DINO SwinB 768 0.79 0.36 0.17 0 0.26 0.05 0 0 0.16 0.08 0.10 0.02 0.60 0.35 0.24 0.04
SwinT 768 0.72 0.42 0.10 0.02 0.25 0.07 0 0 0.08 0.13 0.04 0.08 0.52 0.34 0.24 0.07
Table 2: Comparison of VFMs with various backbones on covariate shift scenarios using GMM and NF methods. Results show FPR95↓ (lower is better) for different environmental conditions: synthetic fog with varying intensities (Foggy Cityscapes [68]), synthetic effects (Bravo-Synflare[50], Bravo-Synrain[45]), and real adverse conditions (ACDC[69]). Best values for each dataset are highlighted in bold.
Architecture-Specific Distribution Modeling: Our analysis reveals that Grounding DINO with Swin Transformer achieves notable performance when paired with flow-based density estimation (Lost and Found: FPR95: 0.16; Bravo-Synobj: FPR95: 0.02). The hierarchical design of Swin Transformer, with its local-global attention mechanism, captures richer semantic relationships than standard architectures. However, leveraging this enhanced semantic understanding requires appropriate distribution modeling—while simpler approaches show limited effectiveness with Swin Transformer (APS, MFS, OC-SVM: FPR95: 0.47–0.54), normalizing flows unlock its full potential. This architecture-specific behavior demonstrates the importance of matching distribution modeling techniques to architectural characteristics.
Real vs. Synthetic Performance: Notably, CLIP’s RN50x64 maintains consistent performance with conventional methods like GMM and OC-SVM on real-world data (FPR95: 0.23–0.24 on Lost and Found), but struggles with synthetic anomalies (FPR95: 0.53–0.54 on Bravo-Synobj). In contrast, Grounding DINO with flow-based modeling maintains effectiveness across both real and synthetic scenarios, suggesting superior generalization to diverse semantic shift types regardless of data origin.
4.2 Covariate Shift
As previously, due to space limitations we present results for only the two top-performing backbones per model in terms of FPR95 scores, focusing on GMM and NF methods due to their superior detection capabilities and FPR95’s sensitivity in distinguishing architectural performance. A comprehensive comparison is available in the Appendix. The following key observations can be drawn from our experiments.
Perturbation Intensity: Performance generally improves with increased perturbation intensity, as evidenced by the Foggy Cityscapes [68] results (i=0.05 vs. i=0.2), indicating sensitivity to the magnitude of covariate shift.
Architecture and Detection Performance: Grounding DINO with Swin Transformer as highly effective in detecting most covariate shifts when paired with normalizing flows, achieving perfect detection (FPR95: 0) on several conditions. However, its performance exhibits an interesting pattern in rain detection: perfect detection on Bravo-Synrain, which features explicit rain drops on the camera lens, but degraded performance on ACDC Rain (FPR95: 0.34-0.35), which only captures environmental rain effects (see example images in Appendix). Similarly, CLIP’s RN50x64 shows higher false positive rates on synthetic perturbations (fog: 0.67, flare: 0.47) but excels in real adverse conditions (ACDC Snow: 0.03), suggesting that architectural designs may be more sensitive to direct camera-level perturbations than to subtle environmental changes.
Model Backbone SegmentMeIfYouCan Indian Driving Dataset
FPR95↓ FPR95↓
GMM NF GMM NF
CLIP RN50x64 0 0 0 0
ViT-B/16 0 0 0 0
DINO ViT-B/16 0.01 0.21 0.26 0.26
ViT-B/8 0 0 0.12 0.07
DINOv2 ViT-L/14 0 0.01 0 0
ViT-g/14 0 0 0 0
Grounding DINO SwinB 0 0 0.01 0
SwinT 0 0 0.05 0
Table 3: Sensitivity of VFMs on datasets combining semantic and covariate shifts. Results are shown for the SMIYC dataset [11] (obstacle track) and IDD [77], using the two top-performing two backbones for each model. Highest scores are highlighted in bold.
4.3 Semantic and Covariate Shift
To evaluate the robustness of VFMs to simultaneous semantic and covariate shifts, we utilized the SMIYC dataset [11] and the Indian Driving Dataset (IDD) [77]. SMIYC introduces both novel objects and adverse weather conditions, while IDD presents a geographically distinct urban environment with unique objects and weather patterns. SMIYC contains two datasets, RoadAnomaly21 and RoadObstacle21, the latter of which was used for results in Tab. 3. As in previous experiments, Cityscapes [16] training images served as the reference, with validation images as ID samples. Tab. 3 shows GMM results for each the two top-performing backbones, the full comparison can be found in the Appendix.
VFMs demonstrate strong robustness to both semantic and covariate shifts, with most models achieving near-perfect FPR95 scores on both the SMIYC and Indian Driving Dataset. This is to be expected as both shifts at a time constitute a stronger deviation from the training data than previously.
4.4 Benchmarking Against OOD Methods and ImageNet-Trained Counterparts
We evaluated our approach against both established OOD detection methods [3] and ImageNet-trained counterparts of the same architectures in the context of input monitoring for AD systems. To facilitate comparison, we implemented these methods using a DeepLabV3+ semantic segmentation network with ResNet50 backbone [^6] DeepLabv3+-RN50 trained on Cityscapes:: https://github.com/open-mmlab/mmsegmentation/tree/main/configs/deeplabv3plus trained on Cityscapes [16] dataset. Our VFM-based input monitoring technique was benchmarked against methods categorized by their detection mechanisms: feature-based approaches (GradNorm [32]), which analyze network activations and gradients; logit-based methods (Energy [46], Entropy [51], MaxLogit [89]), which examine output characteristics; confidence-based methods (DICE [75]), which leverage confidence scores derived from the model; distance-based approaches (KNN [76]), which assess feature space similarities; and ODIN [42], which enhances OOD detection through temperature scaling and input preprocessing. These methods require specific network access: intermediate layers for feature-based, final layer for logit-based, and feature embeddings for distance-based approaches. In contrast, our VFM-based method operates independently of the monitored network, significantly enhancing its applicability in AD systems where network architectures may be proprietary or inaccessible.
For our experimental evaluation, we used Cityscapes [16] training set as $\mathcal{X}\mathrm{monitor}$ and its validation set as $\mathcal{X}_{\mathrm{ID}}$, testing across diverse OOD scenarios: real-world obstacles (Lost and Found [59]), synthetic anomalies (Bravo-SynObj [50]), adverse weather conditions (Foggy Cityscapes [68], Bravo-SynRain [50]). For fair comparison, all methods were evaluated using ResNet50 backbones from both the segmentation network and CLIP, with our approach utilizing GMM and NF for distribution modeling. Additionally, we report results using Grounding DINO’s SwinB backbone with NF, which achieved the best performance overall. The results are shown in Tab. 4.
Method Lost and Found Bravo-Synobj Foggy Cityscapes (i=0.2) Bravo-Synrain
AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95
OOD Methods DICE [75] 0.50 0.94 0.57 0.91 0.69 0.92 0.77 0.80
Energy [46] 0.50 0.97 0.50 0.97 0.49 0.99 0.50 0.99
Entropy [51] 0.60 0.88 0.61 0.86 0.64 0.90 0.70 0.86
GradNorm [32] 0.80 0.43 0.45 0.94 0.98 0.06 0.99 0.00
KNN [76] 0.68 0.78 0.49 0.91 0.57 0.88 0.42 0.94
MaxLogit [89] 0.56 0.60 0.56 0.95 0.63 0.95 0.62 0.97
ODIN [42] 0.56 0.94 0.56 0.95 0.59 0.93 0.58 0.97
Image-Net Trained Counterparts RN50-GMM 0.76 0.62 0.84 0.75 0.92 0.28 0.94 0.23
RN50-NF 0.75 0.63 0.87 0.74 0.93 0.25 0.97 0.12
SwinB-NF 0.86 0.44 0.91 0.53 0.92 0.35 0.99 0.05
VFMs Ours (RN50-GMM) 0.90 0.36 0.89 0.66 0.85 0.47 0.85 0.43
Ours (RN50-NF) 0.82 0.52 0.92 0.57 0.86 0.53 0.95 0.16
Ours (SwinB-NF) 0.95 0.16 0.99 0.02 1.00 0.00 1.00 0.02
Table 4: Comparison of OOD detection between baseline methods and our VFM-based approach across semantic shifts (Lost and Found, Bravo-Synobj) and covariate shifts (Foggy Cityscapes, Bravo-Synrain). All baselines use a DeepLabV3+ network with ResNet50 backbone. Our approach is evaluated with CLIP’s RN50 using GMM and NF for distribution modeling. Best results per dataset in bold.
OOD detection methods show varying performance across semantic and covariate shifts. In particular, confidence-based methods (DICE [75]) and distance-based methods (KNN [76]) demonstrate moderate performance on Lost and Found (AUROC: 0.50-0.68) and Bravo-Synobj (AUROC: 0.49-0.57), suggesting their limited capability in capturing complex semantic shifts. This limitation likely arises from their reliance on model confidence scores (DICE) and feature similarity measures (KNN), which are often unreliable under high semantic variations.
Logit-based approaches (Energy [46], Entropy [51], MaxLogit [89]) perform consistently but modestly across all datasets (AUROC: 0.50-0.70), indicating their relative robustness but limited discriminative power. Their uniform performance suggests that while these methods leverage logit magnitudes to separate ID and OOD samples, they may struggle with fine-grained anomalies that do not cause strong logit deviations.
Notably, feature-based methods like GradNorm [32] show strong performance on covariate shifts (Foggy Cityscapes AUROC: 0.98, FPR95: 0.06; Bravo-Synrain AUROC: 0.99, FPR95: 0.00) and perform reasonably well on real semantic shifts (Lost and Found AUROC: 0.80, FPR95: 0.43) but struggle significantly with synthetic anomalies (Bravo-Synobj AUROC: 0.45, FPR95: 0.94), likely because GradNorm is more sensitive to global image transformations and natural distortions, while synthetic object insertions do not trigger strong gradient deviations.
ImageNet-trained models perform well on covariate shifts but lag behind VFMs on semantic shifts. Using identical architectures (ResNet50, SwinB) trained solely on ImageNet with the same distribution modeling techniques, we found competitive performance on covariate shifts. SwinB-NF excelled on weather conditions (Bravo-Synrain: AUROC: 0.99, FPR95: 0.05) and RN50-NF performed strongly on fog detection (Foggy Cityscapes: AUROC: 0.93, FPR95: 0.25). However, a significant gap appears for semantic shifts, particularly with SwinB (ImageNet SwinB-NF on Bravo-Synobj: FPR95: 0.53 vs. VFM SwinB-NF: FPR95: 0.02). This suggests ImageNet pre-training captures features sufficient for environmental variations, while VFMs’ diverse training objectives and broader data exposure significantly enhance detection of novel objects and semantic anomalies.
4.5 Input Monitoring
To evaluate the practical applicability of our input monitoring approach, we applied it to semantic segmentation using DeepLabv3+ [14] with ResNet-50, ResNet-101 [26], and MobileNetV3 [31] backbones. These models were trained on Cityscapes [16] and evaluated on IDD [77]. We employed Grounding DINO’s SwinB backbone with normalizing flows to model the distribution of Cityscapes training images. For each test image in IDD, we computed its log-likelihood under this learned distribution, providing a measure of similarity to the training distribution. To systematically evaluate the effect of input monitoring, we established baselines using the entire IDD test set, then applied increasingly selective filtering based on the computed log-likelihoods. Specifically, we defined three filtering thresholds that retain the top 75% (low), 50% (medium), and 25% (high) of IDD images ranked by their log-likelihood scores. Segmentation performance was evaluated using mean Intersection over Union (mIoU) and mean Accuracy (mAcc) for each filtering level.
Filter Strength ResNet50 ResNet101 MobileNetv3
mIoU↑ mAcc↑ mIoU↑ mAcc↑ mIoU↑ mAcc↑
None 43.58 62.84 45.85 65.1 34.22 52.25
Low 44.73 63.62 46.81 65.81 35.4 53.32
Medium 46.86 64.82 48.35 66.54 36.59 54.11
High 49.38 65.19 50.52 66.42 38.36 54.02
Table 5: Semantic segmentation performance of DeepLabv3+ with ResNet-50, ResNet-101, and MobileNetV3 backbones on IDD for different input monitoring filter strengths.
Results are summarized in Tab. 5. Our input monitoring approach clearly enhances semantic segmentation performance on OOD data. Utilizing Grounding DINO’s [45] SwinB backbone with NF, the method quantifies distributional shifts between test and training sets, enabling adaptive input filtering. The results reveal a clear trade-off: stricter monitoring yields fewer retained images but clear accuracy gains. Stronger performance improvements can be expected on even more heterogeneous datasets, containing, e.g., Indian and German driving scenes.
5 Conclusion
We presented an unsupervised, model-agnostic binary OOD classification framework that leverages latent representations of pre-trained off-the-shelf VFMs together with density modeling for joint detection of both semantic and covariate shifts. This is for the first time thoroughly tested in the challenging domain of AD. Our extensive experiments show that simple techniques for modeling the VFM feature distributions are sufficient to detect semantic and/or covariate OOD shifts with astonishingly high accuracy. VFMs clearly outperform non-foundation model latent representations, and classical OOD methods. Lastly, despite our method being model-agnostic, down-stream perception modules exhibit clear performance improvements when filtering out data that deviates too much from the training distribution. This promises that VMF latent space distributions can help to identify inputs with higher risk of inducing errors in computer vision DNNs, even in complex open-world domains.
Acknowledgment
We acknowledge the financial support by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) within the project nxtAIM – Next Generation AI Methods and the German Federal Ministry of Education and Research (BMBF) within the project REFRAME (grant no. 01IS24073).
Supplementary Material
This supplementary material provides comprehensive analyses and additional details that complement the main paper. Concretely, details are provided on:
  • A: reasoning for the scope
  • B: model and dataset properties with particular emphasis on both semantic and covariate distribution shift coverage, supported by illustrative examples.
  • C: implementation considerations essential for reproducing our experimental framework.
  • D: additional tables with complete experimental results across all model architectures and backbones
  • E: More detailed discussion of limitations and future directions.
Appendix A Discussion of the Scope
This section further argues specific choices of scope:
  • A.1: Why input monitoring (Comparison with pixel-wise OOD segmentation)
  • A.2: Why VFMs instead of standard feature extractors
  • A.3 Why real-time capability may reasonably be achieved
A.1 Input Monitoring and Pixel-Level OOD Segmentation
Lost and Found Dataset
ACDC Night Dataset
ACDC Rain Dataset
SegmentMeIfYouCan Anomaly Track Dataset
Figure 2: Visualization of Mask2Anomaly [63] applied to selected images exhibiting semantic and covariate shifts. The left column presents the original images from various datasets, including Lost and Found [59], ACDC Night and Rain [69], and SegmentMeIfYouCan Anomaly Track [11] The right column displays the corresponding OOD object-level maps generated by Mask2Anomaly.
Pixel-level out-of-distribution (OOD) segmentation methods aim to identify pixels that deviate from the known distribution of semantic classes in training data. This method is about image-level OOD detection (“is there any anomaly”), not pixel-level OOD segmentation (“where is an anomaly”). We detail the motivation for this in the following.
Practically, input monitoring complements pixel-wise OOD segmentation by enabling more efficient anomaly detection and response for the following reasons:
  • In case there is a shift in one part of the image, the probability that there are further (undetected) shifts elsewhere in the image is very high. Therefore, a reprocessing of the complete image using a safety mechanism is recommended. This requires an accurate prediction of whether there is any anomaly in the image. And despite considerable progress in pixel-level OOD segmentation [4, 24, 78, 65, 53, 12, 37, 63, 24, 87, 23, 47], these are still challenged by a high rate of false positives (illustrated in Fig. 2). This lies in the nature of problem formulation: Ensuring that there is no false positive among the tremendous amount of pixels in an image requires way more accurate predictions than a single image level OOD detection. This issue is illustrated in Fig. 2 via representative examples of Mask2Anomaly [63], a leading method from the SMIYC benchmark [11]. While pixel-wise OOD segmentation provides detailed information about environmental anomalies, its practical deployment in safety-critical applications remains challenged by high false positive rates.
  • Furthermore, pixel-level OOD detection methods typically require supervision during training and exhibit performance degradation when encountering objects or scenes outside their training domain. Consequently, deploying pixel-level OOD segmentation methods in real-world applications requires an additional system capable of determining when these models can be reliably employed. Given that the operational domain of AD encompasses diverse semantic and covariate shifts, deploying such methods necessitates an additional system to determine their applicability.
These limitations underscore the importance of operation-level input monitoring for robust assessment of perception system reliability across varying environmental conditions.
Recently, Gao et al. [22] addressed the challenge of pixel-level semantic segmentation under domain shift by leveraging generative models for training data augmentation. While their implementation is not yet publicly available, evaluating their approach across diverse covariate shift scenarios presents a promising direction for future investigation.
A.2 Why VFMs for Feature Extraction Instead of Standard Vision Models
In this section we provide evidence that VFM feature encodings are clearly superior to those of classical feature extractor DNNs. This motivated our choice to restrict further analysis of OOD detection capabilities to the scope of VFM feature encodings.
To contextualize the performance of VFMs, we establish comparative baselines using analogous architectures pretrained on ImageNet [18] and Autoencoders (AEs) trained on our target domain data. These experiments provide detailed evidence supporting our main paper’s finding that VFMs outperform non-foundation model latent representations in input monitoring. Our baseline models include ImageNet-pretrained architectures: ResNet variants (e.g. RN50, RN101, RN152) [26], Vision Transformers (e.g. ViT-B/16, ViT-L/16) [20], and Swin Transformers (i.e. SwinB, SwinT) [48] [^7] Implementation details: ResNet models from torchvision (https://github.com/pytorch/vision), ViT and Swin Transformer models from timm (https://github.com/huggingface/pytorch-image-models). Additionally, we train domain-specific AEs with varying latent space dimensions on the Cityscapes training set [16].
Baseline Setup
For all architectures (ImageNet-pretrained models and AEs), we model their latent space distributions using the approaches presented in the main paper: APS, OC-SVM, GMM, and NF. We evaluate these models using the Cityscapes dataset, where the training set serves as $\mathcal{X}_{\mathrm{monitor}}$ and the validation set as $\mathcal{X}_{\mathrm{ID}}$. To ensure comprehensive evaluation, we test on both semantic and covariate shift scenarios. Detailed implementation specifications and architectural configurations are provided in Appendix C.
Results
Based on the comparative analysis in Tab. 6, VFMs consistently outperform ImageNet-trained models and AEs across datasets and different shifts, i.e. semantic and covariate. We observe three critical findings through the lens of FPR95: First, VFMs demonstrate superior false positive control, with CLIP-RN50x64 achieving FPR95 of 0.24 on Lost and Found, outperforming ImageNet-trained counterparts. Second, the choice of density estimator proves crucial - while GMM provides competitive results, normalizing flows unlock exceptional performance for hierarchical architectures, enabling Grounding DINO’s SwinB to achieve state-of-the-art FPR95 of 0.16 on Lost and Found and perfect detection (FPR95 = 0.00) on Bravo-Synrain. Finally, despite their structural simplicity, autoencoders struggle with false positive control (FPR95 $>$ 0.65 across all configurations), highlighting that architectural sophistication in VFMs, rather than mere representation learning, is key to reliable input monitoring.
Model Backbone Lost and Found
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
Trained on ImageNet RN50 0.73 0.69 0.70 0.73 0.69 0.70 0.76 0.71 0.62 0.70 0.67 0.78 0.75 0.71 0.63
RN101 0.70 0.64 0.68 0.70 0.64 0.68 0.72 0.65 0.60 0.69 0.65 0.81 0.75 0.70 0.63
RN152 0.68 0.63 0.74 0.68 0.63 0.74 0.72 0.68 0.77 0.67 0.62 0.77 0.70 0.65 0.73
ViT-B/16 0.79 0.75 0.53 0.79 0.75 0.53 0.83 0.81 0.53 0.81 0.78 0.53 0.82 0.79 0.56
ViT-L/16 0.88 0.85 0.39 0.88 0.85 0.39 0.86 0.83 0.47 0.88 0.85 0.40 0.88 0.85 0.42
SwinB 0.84 0.82 0.53 0.84 0.82 0.53 0.83 0.84 0.54 0.85 0.84 0.53 0.86 0.85 0.44
SwinT 0.72 0.64 0.62 0.72 0.64 0.62 0.74 0.70 0.75 0.73 0.66 0.65 0.72 0.67 0.77
Autoencoder 128 0.66 0.61 0.71 0.55 0.52 0.80 0.62 0.56 0.74 0.62 0.55 0.79 0.42 0.43 0.94
256 0.68 0.62 0.69 0.70 0.67 0.69 0.68 0.64 0.77 0.52 0.48 0.87 0.60 0.56 0.82
512 0.65 0.61 0.73 0.68 0.63 0.70 0.62 0.58 0.74 0.61 0.58 0.88 0.55 0.53 0.88
1024 0.70 0.67 0.74 0.77 0.74 0.56 0.62 0.57 0.84 0.65 0.60 0.83 0.68 0.63 0.65
Ours Clip-RN50 0.85 0.82 0.49 0.85 0.82 0.49 0.90 0.88 0.36 0.85 0.83 0.51 0.82 0.78 0.52
Clip-RN50x64 0.94 0.93 0.24 0.94 0.93 0.24 0.94 0.93 0.23 0.94 0.93 0.24 0.88 0.84 0.37
SwinB 0.88 0.84 0.47 0.88 0.84 0.47 0.91 0.89 0.32 0.84 0.80 0.54 0.95 0.94 0.16
Bravo Synrain
Trained on ImageNet RN50 0.90 0.89 0.33 0.79 0.83 0.82 0.94 0.93 0.23 0.92 0.91 0.27 0.97 0.97 0.12
RN101 0.89 0.87 0.36 0.81 0.85 0.79 0.93 0.92 0.27 0.91 0.90 0.30 0.97 0.96 0.12
RN152 0.90 0.89 0.35 0.80 0.84 0.78 0.95 0.94 0.18 0.93 0.91 0.29 0.97 0.96 0.10
ViT-B/16 0.86 0.82 0.38 0.81 0.81 0.69 0.93 0.92 0.24 0.91 0.88 0.26 0.97 0.96 0.12
ViT-L/16 0.88 0.84 0.31 0.85 0.88 0.70 0.92 0.91 0.24 0.91 0.88 0.23 0.98 0.97 0.07
SwinB 0.90 0.88 0.31 0.84 0.86 0.67 0.96 0.96 0.14 0.95 0.93 0.19 0.99 0.99 0.05
SwinT 0.88 0.84 0.31 0.76 0.75 0.77 0.95 0.93 0.16 0.93 0.90 0.20 0.99 0.98 0.04
Autoencoder 128 0.60 0.59 0.87 0.58 0.55 0.87 0.59 0.56 0.86 0.60 0.58 0.88 0.55 0.52 0.87
256 0.60 0.59 0.85 0.57 0.55 0.87 0.58 0.55 0.87 0.54 0.53 0.90 0.59 0.55 0.85
512 0.60 0.60 0.84 0.56 0.54 0.88 0.56 0.55 0.87 0.60 0.59 0.88 0.57 0.54 0.85
1024 0.57 0.55 0.87 0.59 0.56 0.86 0.60 0.57 0.86 0.52 0.51 0.86 0.63 0.59 0.81
Ours Clip-RN50 0.76 0.72 0.59 0.76 0.72 0.59 0.85 0.82 0.43 0.75 0.71 0.64 0.95 0.92 0.16
Clip-RN50x64 0.99 0.99 0.03 0.99 0.99 0.03 1.00 1.00 0.01 1.00 1.00 0.01 0.99 0.99 0.04
SwinB 1.00 1.00 0.02 1.00 1.00 0.02 1.00 1.00 0.00 1.00 0.99 0.02 1.00 1.00 0.00
Table 6: Comparative evaluation of out-of-distribution detection performance using ImageNet-pretrained models, autoencoders, and VFMs on Lost and Found [59] and Bravo-Synrain [50] benchmarks. We evaluate using AUROC, AUPR, and FPR95 metrics across five methods: APS, MFS, OC-SVM, GMM, and NF. For VFMs, we utilize image encoders from CLIP [61] with ResNet-50 and ResNet-50x64 backbones, and from Grounding-DINO [45] with SwinB backbone. Best performances for each metric and dataset are highlighted in bold.
A.3 Real-time Capability of Chosen Models
A critical consideration for practical deployment in AD systems is computational efficiency. Our approach leverages only the image encoder components of Vision Foundation Models (VFMs), specifically employing CLIP’s ResNet variants [61] and Grounding DINO’s Swin Transformer backbones [45]. Recent literature substantiates the real-time capabilities of these architectures:
  • Zhao et al.[90] demonstrate that DETR-based architectures achieve real-time performance while surpassing YOLO-based models in detection accuracy.
  • Kassab et al.[36] have successfully deployed CLIP’s ResNet backbones for real-time visual scene understanding in SLAM applications, validating their computational feasibility.
Our framework offers additional deployment flexibility through temporal and spatial optimization strategies, including:
  • Processing every $n$th frame to reduce computational load
  • Operating on downsampled inputs while maintaining monitoring efficacy
To quantify real-time performance, we benchmarked all feature extractors on hardware (Intel Xeon Gold 6326 CPU, NVIDIA A40 GPU) using randomly sampled 50 full-resolution (2048×1024) images from Cityscapes. Table 7 presents frames-per-second (FPS) rates with standard deviations across 5 measurement runs.
Model Backbone FPS
Mean Std
CLIP RN50 63.69 0.81
RN101 63.29 0.40
RN50x4 56.50 0.32
RN50x16 42.37 0.90
RN50x64 31.15 0.19
ViT-B/32 63.29 1.60
ViT-B/16 56.18 0.32
ViT-L/14 49.50 0.74
DINO ViT-S/16 59.17 2.10
ViT-S/8 49.50 0.00
ViT-B/16 49.02 3.37
ViT-B/8 42.74 0.18
RN50 56.50 0.00
DINOv2 ViT-S/14 53.19 1.70
ViT-B/14 57.14 2.94
ViT-L/14 47.62 0.23
ViT-G/14 35.34 0.87
Grounding DINO SwinB 16.31 0.08
SwinT 26.74 2.43
Table 7: Inference speed of VFMs’ feature extractors, measured in frames per second (FPS) per image on high-resolution 2048×1024 input images. The FPS values represent the mean and standard deviation (Std) across multiple runs.
Our results indicate that CLIP’s ResNet and ViT backbones provide the highest throughput, with RN50 and ViT-B/32 both exceeding 63 FPS. The DINOv2 and DINO variants maintain strong performance between 35-59 FPS. While Grounding DINO models demonstrate lower throughput (SwinB: 16.31 FPS, SwinT: 26.74 FPS), they remain viable for real-time applications—particularly given their superior OOD detection performance. These throughput rates can be improved by employing our proposed temporal optimization (processing every $n$th frame) or spatial optimization (downsampling input resolution). Recent work by Zhao et al. [90] has shown that DETR-based architectures can surpass YOLO [15] in the inference with their optimizations which can be used in our case for Grounding-DINO measurements.
For distribution modeling approaches, computational efficiency varies significantly. APS requires computing pairwise similarities between all monitor and test samples, making it computationally prohibitive for real-time applications. In contrast, MFS, OC-SVM, GMM, and NF all offer efficient inference after offline training, as they only require evaluating new samples against pre-computed distribution parameters. Among these approaches, MFS provides the fastest inference but with reduced detection performance, while GMM and NF offer optimal performance-efficiency trade-offs for operational deployment.
Appendix B Details of Chosen Datasets and Vision Foundation Models
B.1 Diversity of Chosen VMFs
In our experiments, we employ four VFMs: CLIP [61], DINO [10], DINOv2 [56], and Grounding DINO [45]. A critical consideration in evaluating these models is whether their superior feature representations might be attributed to test data leakage. As shown below, the chosen models
  • cover a wide range of architectures, as well as
  • a wide range of different training datasets; and
  • no test data leakage influenced our results.
CLIP [61]
was trained on a proprietary dataset of 400 million image-text pairs curated from publicly available internet data, leveraging natural language descriptions paired with corresponding images. In contrast, our datasets—e.g. Cityscapes [16], ACDC [69], and Lost and Found [59]—are specifically designed for perception tasks, offering pixel-level semantic labels rather than natural language annotations. Moreover, these datasets’ highly specialized characteristics—featuring calibrated front-facing cameras, standardized viewpoints, and tightly controlled acquisition protocols—differ fundamentally from the diverse, unconstrained nature of web-scraped imagery.
DINO and DINOv2
employ self-supervised learning paradigms, representing a distinct approach to training. While DINO is trained exclusively on ImageNet [66] using self-distillation without labeled data, DINOv2 leverages the larger LVD-142M dataset [56]. Although Cityscapes and ACDC are part of the LVD-142M corpus, our strongest results are achieved using CLIP and Grounding-DINO rather than DINOv2, suggesting that performance stems from architectural strengths rather than data memorization.
Grounding-DINO [45]
demonstrates strong performance in our experiments, and was trained on detection datasets (COCO [43], Objects365 [72], OpenImage [39]) and grounding datasets grouped as GoldG [35] (Flickr30k entities [60], Visual Genome [40]) and RefC (RefCOCO variants)[35]. Notably, our evaluation datasets were not part of its training corpus.
These findings indicate that VFMs’ superior performance stems from their architectural and representational capabilities rather than training data overlap or test data leakage.
B.2 Dataset Examples and Characteristics
This section provides a comprehensive overview (see also Tab. 8) and samples of the datasets employed in our experimental framework and their strategic utilization for evaluating different types of distribution shifts. We categorize our evaluation scenarios into three distinct types of shifts:
  1. 1.
    semantic shifts (novel objects or classes),
  2. 2.
    covariate shifts (environmental and capture condition variations), and
  3. 3.
    combined shifts that exhibit both semantic and covariate variations simultaneously.
(a) ACDC Fog
(b) ACDC Night
(c) ACDC Rain
(d) ACDC Snow
Figure 3: Example images from four subsets of ACDC Dataset [69]
ACDC Dataset:
The ACDC dataset [69] is designed for semantic scene understanding under adverse driving conditions, comprising four distinct weather scenarios: fog, night, rain, and snow. The dataset contains 4006 images with fine-grained pixel-level panoptic annotations, where each adverse condition image is paired with a corresponding reference image captured under clear conditions at the same location. Example images from each weather scenario are shown in Figure 3.
For our experiments, we leverage the dataset’s paired structure: reference images from the training set serve as $\mathcal{X}_{\mathrm{monitor}}$, reference images from the validation set as $\mathcal{X}_{\mathrm{ID}}$, and images from each adverse condition (i.e. fog, night, rain, and snow) as separate $\mathcal{X}_{\mathrm{test}}$ sets. This setup enables systematic evaluation of covariate shift detection across different environmental conditions while maintaining consistent scene content.
(a) Bravo-Synflare
(b) Bravo-Synobjs
(c) Bravo-Synrain
(d) Lost and Found
Figure 4: Example images from Bravo datasets [50] and Lost and Found Dataset [59]
Bravo Datasets:
The Bravo dataset family [50] provides synthetic perturbations of Cityscapes images to evaluate model robustness under controlled distribution shifts. The collection comprises three distinct subsets:
  • Bravo-Synflare (lens flare effects simulating intense light sources),
  • Bravo-Synobjs (synthetic object insertions for semantic shifts), and
  • Bravo-Synrain (synthetic precipitation effects).
Examples from each subset are shown in Fig. 4.
We structure our experiments using Cityscapes training data as $\mathcal{X}_{\mathrm{monitor}}$ and validation data as $\mathcal{X}_{\mathrm{ID}}$. For $\mathcal{X}_{\mathrm{test}}$, we evaluate covariate shifts using Bravo-Synflare and Bravo-Synrain, while semantic shifts are assessed using Bravo-Synobjs. This configuration enables systematic evaluation of both types of distribution shifts while maintaining consistent base scene content from the Cityscapes domain.
Lost and Found Dataset:
The Lost and Found dataset [59] is designed for real-world anomaly detection in urban driving environments, specifically focusing on rare and unexpected road hazards such as traffic cones, cargo items, and other small obstacles that are typically absent from standard training datasets. Captured in German cities under similar environmental conditions as Cityscapes [16], it maintains consistent scene characteristics while introducing novel objects, making it particularly suitable for evaluating semantic shifts in the Cityscapes domain. Example images from this dataset are shown in Fig. 4.
For our experiments, we utilize Cityscapes training data as $\mathcal{X}_{\mathrm{monitor}}$ and its validation set as $\mathcal{X}_{\mathrm{ID}}$, while Lost and Found samples serve as $\mathcal{X}_{\mathrm{test}}$. This setup enables evaluation of semantic shift detection capabilities in naturally occurring, safety-critical scenarios while minimizing covariate shifts due to shared geographical and capture conditions.
(a) Cityscapes
(b) Foggy Cityscapes
(c) Indian Driving Dataset
(d) SegmentMeIfYouCan
Figure 5: Example images from Cityscapes [16], Cityscapes Fog [68], Indian Driving dataset[77] and SegmentMeIfYouCan [11]
Cityscapes and Foggy Cityscapes Datasets:
The Cityscapes [16] dataset is a widely-used benchmark for semantic urban scene understanding. It contains high-quality pixel-level annotations of images from 50 cities across Germany, captured in various urban driving conditions. The Foggy Cityscapes Dataset is an adaptation of the original Cityscapes dataset, created by synthetically adding fog to the images. Example images can be seen in Figure Fig. 5.
Indian Driving Dataset:
The Indian Driving dataset [77] (IDD) is a real-world dataset focused on capturing the complexities of driving scenes in Indian road conditions. It consists of images collected from various urban, rural, and highway scenes across India, representing a diverse set of traffic conditions. An example image can be seen in Figure Fig. 5.
SegmentMeIfYouCan Dataset:
The SegmentMeIfYouCan [11] dataset (SMYIC) is designed for evaluating anomaly detection in semantic segmentation tasks. It includes challenging scenarios with diverse visual contexts and synthetic anomalies that closely resemble real-world driving conditions. The dataset focuses on identifying OOD objects that were not present during training. An example image can be seen in Figure Fig. 5.
In our semantic and covariate shift experiments presented in the main paper, we maintain Cityscapes training data as $\mathcal{X}_{\mathrm{monitor}}$ and its validation set as $\mathcal{X}_{\mathrm{ID}}$, while samples from IDD [77] and SMIYC [11] serve as $\mathcal{X}_{\mathrm{test}}$. This setup enables evaluation of our approach under real-world scenarios where both types of shifts occur simultaneously.
ID monitor train data $\mathcal{X}_{\text{monitor}}$ ID test data $\mathcal{X}_{\text{ID}}$ OOD test data $\mathcal{X}_{\text{test}}$
Semantic Shift Covariate Shift Semantic and Covariate Shift
Cityscapes
train split
Cityscapes
validation split
Lost and Found
Bravo-Synobj
Foggy Cityscapes
Bravo-Synflare
Bravo-Synrain
Indian Driving Dataset
SegmentMeIfYouCan
ACDC
(clear weather)
train split
ACDC
(clear weather)
validation split
- ACDC-Fog
ACDC-Night
ACDC-Rain
ACDC-Snow
Table 8: Dataset organization for evaluating different types of distribution shifts.
Table 8 illustrates our systematic dataset organization. For Cityscapes, we utilize its training set as $\mathcal{X}_{\mathrm{monitor}}$ and validation set as $\mathcal{X}_{\mathrm{ID}}$. Similarly for ACDC, we use the reference (clear-weather) images from its training set as $\mathcal{X}_{\mathrm{monitor}}$ and reference validation set as $\mathcal{X}_{\mathrm{ID}}$. The remaining datasets are categorized into three distinct evaluation scenarios: semantic shifts (novel objects), covariate shifts (environmental variations), and combined semantic and covariate shifts, each serving as $\mathcal{X}_{\mathrm{test}}$ for their respective experiments.
Appendix C Implementation Details
In this section, we detail our feature processing pipeline and implementation specifics. First, we describe the flattening procedure for converting 3-dimensional features from Grounding DINO’s [45] image encoder into single-dimensional vectors. We then present the architecture design and training protocol for the autoencoder models evaluated in Sec. A.2. Finally, we outline our systematic parameter selection process, including kernel optimization for each backbone architecture and the hyperparameter configuration of our normalizing flow models.
Figure 6: Comparison of AIC values across various model backbones, highlighting the trade-off between model complexity and goodness of fit for GMMs applied to different VFMs. Each plot displays the AIC values for a specific backbone architecture, with the optimal number of GMM components (minimizing AIC) indicated. The size of the latent space vector for each architecture is also annotated.
Selecting Number of Components for GMMs
To determine the optimal number of components / Gaussians $K$ for the GMM used to model the feature distributions, we employ the Akaike Information Criterion (AIC). The AIC is a model selection technique that strikes a balance between the goodness of fit and the complexity of the model, thereby preventing overfitting. The AIC for a given GMM with $K$ components is defined as
$\text{AIC}(K)=2k(K)-2\ln(L)$ (10)
where $k$ is the number of parameters in the model, and $L$ is the value of the likelihood function for the estimated model, i.e., after likelihood maximization. For a GMM with $K$ components, each with $d$-dimensional mean vectors and full covariance matrices, the number of parameters is
$k(K)=K\left(d+\frac{1}{2}d(d+1)\right)-1$ (11)
The lower the AIC value, the better the trade-off between goodness of fit and model complexity. To select the optimal $K$, we fit GMMs with varying numbers of components (e.g., $K=1,2,3,\ldots$) to a hold-out validation set representative of the in-distribution data. For each value of $K$, we compute the corresponding $\text{AIC}(K)$. We then select the value of $K$ that minimizes the AIC, as this represents a balance between model fit and complexity for the given data, which in theory can be optimal under certain assumptions.
Figure 6 presents a comprehensive evaluation of VFMs. Additionally, Figure 7 illustrates the AIC values for ImageNet-trained backbones and autoencoders trained on reference data, corresponding to the experimental setup detailed in Sec. A.2. Each subplot within these figures represents the AIC profile for a specific backbone architecture, elucidating the intricate trade-off between model complexity (quantified by the number of GMM components) and goodness of fit. The optimal number of components, denoted as $K$, is identified for each backbone as the value corresponding to the global minimum of the AIC curve. This optimal point is explicitly marked on each graph, facilitating direct comparison across different architectural choices. This systematic approach to model selection provides us with a model complexity of the GMM appropriately tailored to the characteristics of each backbone’s feature space, potentially enhancing the robustness and efficiency of subsequent anomaly detection tasks.
Figure 7: AIC values for ImageNet-trained backbones and autoencoders trained on reference data. Each subplot depicts the AIC profile for a specific architecture or autoencoder configuration, demonstrating the relationship between GMM complexity and fit. The optimal number of components, minimizing AIC, is indicated on each curve.
Grounding DINO Feature Extraction
For feature extraction, we leverage the image encoder of GroundingDINO [45], which is based on the Swin Transformer [48] architecture. Specifically, we extract features from the last three layers of the Swin Transformer and concatenate them to create a multi-scale representation. This approach is analogous to the multi-scale feature utilization in GroundingDINO [45].
Autoencoder Architecture and Training
Layer
Encoder
Input
Conv2d (3, 32, 3x3, stride=2, padding=1)
BatchNorm2d
LeakyReLU
Conv2d (32, 64, 3x3, stride=2, padding=1)
BatchNorm2d
LeakyReLU
Conv2d (64, 128, 3x3, stride=2, padding=1)
BatchNorm2d
LeakyReLU
Conv2d (128, 256, 3x3, stride=2, padding=1)
BatchNorm2d
LeakyReLU
Flatten
Linear (50,176 → Latent Size)
Decoder
Linear (Latent Size → 50,176)
Unflatten
ConvTranspose2d (256, 128, 3x3, stride=2, padding=1)
BatchNorm2d
LeakyReLU
ConvTranspose2d (128, 64, 3x3, stride=2, padding=1)
BatchNorm2d
LeakyReLU
ConvTranspose2d (64, 32, 3x3, stride=2, padding=1)
BatchNorm2d
LeakyReLU
ConvTranspose2d (32, 3, 3x3, stride=2, padding=1)
Sigmoid
Table 9: Architecture of the Autoencoders
The autoencoder is designed to learn compact, latent representations of the reference input data. Once trained, the autoencoder serves as a feature extractor, providing a comparative baseline to compare with the VFMs. The architecture of the autoencoder is detailed in Tab. 9.
The training objective for the autoencoder is to minimize the reconstruction error using the mean squared error loss function, defined as
$\mathcal{L}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\hat{x}_{i})^{2}$ (12)
where $x_{i}$ represents the input feature vector, and $\hat{x}_{i}$ denotes the corresponding reconstructed output. The model is optimized using the Adam optimizer, with a learning rate of $1\times 10^{-3}$. It is trained for 100 epochs. Early stopping is adopted, triggered by a lack of improvement in the loss.
One-Class SVM Optimization
To optimize the performance of the One-Class SVM [1, 8], we conducted a parameter search across various kernel types and their corresponding hyperparameters. The optimization process focused on three kernel functions: radial basis function (RBF), linear, and polynomial, with the aim of selecting the parameter combination that best fits the ID data. This fit was determined based on the mean decision function score on the training set, with higher scores indicating better alignment with the data.
The parameter space explored in this search includes the following:
  • Kernel types: RBF, linear, polynomial
  • $\nu$ values: 0.01, 0.1, 0.5
  • $\gamma$ values (for RBF kernel): scale, auto, 0.1, 1, 10
  • Degree values (for polynomial kernel): 2, 3
For each backbone type, we systematically evaluated combinations of hyperparameters, selecting the configuration that yielded the highest mean decision function score on the reference data set.
Normalizing Flow Implementation
For density estimation using NF, we implement a Real-valued Non-Volume Preserving (Real-NVP) architecture [19]. Our implementation consists of multiple coupling layers with permutation layers and batch normalization between them. Each coupling layer employs a neural network with two hidden layers and ReLU activations, following a half-masking strategy where the first half of the dimensions are used to compute the scale and translation parameters for the second half. We conduct extensive hyperparameter optimization by exploring different architectures with coupling layer topologies of [64, 64], [128, 128], and [256, 256] hidden units, and varying the number of flow steps (2, 4, or 6). The models are trained using the Adam optimizer with a learning rate of $10^{-4}$ and batch sizes of 16 or 32, for either 100 or 200 epochs, using an 80-20 train-validation split for model selection. Following standard practice in flow-based models, we use a standard normal distribution as the base distribution and compute log-probabilities using the change of variables formula with the Jacobian determinant of the transformation.
Appendix D Detailed Experimental Results
This section presents the complete set of experimental results discussed in the main paper. In particular. we provide additional tables detailing the performance metrics across all model architectures and backbone configurations for both semantic and covariate shift detection tasks (not only the top 2 backbones as in the space constraint presentation in the main paper).
D.1 Semantic Shift Experiments
Model Backbone Lost and Found
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.85 0.82 0.49 0.84 0.82 0.52 0.85 0.83 0.51 0.90 0.88 0.36 0.82 0.78 0.52
RN101 0.83 0.82 0.55 0.82 0.81 0.58 0.83 0.82 0.65 0.87 0.85 0.46 0.79 0.74 0.60
RN50x4 0.84 0.81 0.55 0.83 0.82 0.58 0.84 0.81 0.62 0.88 0.86 0.42 0.84 0.79 0.50
RN50x16 0.88 0.86 0.44 0.88 0.88 0.46 0.89 0.88 0.39 0.89 0.87 0.38 0.83 0.80 0.51
RN50x64 0.94 0.93 0.24 0.93 0.93 0.27 0.94 0.93 0.24 0.94 0.93 0.23 0.88 0.84 0.37
ViT-B/32 0.86 0.84 0.51 0.83 0.82 0.60 0.86 0.85 0.53 0.88 0.85 0.38 0.87 0.84 0.50
ViT-B/16 0.88 0.86 0.38 0.86 0.85 0.41 0.89 0.88 0.43 0.91 0.89 0.33 0.91 0.88 0.34
ViT-L/14 0.87 0.85 0.45 0.89 0.89 0.41 0.89 0.87 0.41 0.85 0.83 0.47 0.88 0.86 0.41
DINO ViT-S/16 0.75 0.71 0.59 0.71 0.63 0.60 0.74 0.71 0.76 0.81 0.78 0.50 0.81 0.78 0.55
ViT-S/8 0.71 0.65 0.61 0.66 0.59 0.66 0.75 0.71 0.60 0.82 0.81 0.51 0.79 0.77 0.60
ViT-B/16 0.79 0.75 0.52 0.72 0.64 0.60 0.79 0.76 0.59 0.84 0.82 0.55 0.75 0.71 0.66
ViT-B/8 0.81 0.76 0.49 0.73 0.64 0.59 0.77 0.75 0.66 0.87 0.86 0.45 0.82 0.81 0.59
RN50 0.64 0.63 0.87 0.65 0.62 0.83 0.61 0.60 0.91 0.68 0.67 0.80 0.61 0.60 0.88
DINOv2 ViT-S/14 0.86 0.80 0.34 0.86 0.80 0.36 0.88 0.83 0.31 0.81 0.76 0.51 0.78 0.74 0.65
ViT-B/14 0.86 0.82 0.45 0.86 0.82 0.44 0.87 0.84 0.53 0.82 0.78 0.54 0.85 0.80 0.48
ViT-L/14 0.89 0.87 0.43 0.91 0.88 0.37 0.84 0.82 0.68 0.82 0.77 0.52 0.83 0.78 0.54
ViT-G/14 0.90 0.87 0.30 0.89 0.86 0.35 0.85 0.82 0.63 0.86 0.78 0.39 0.83 0.75 0.46
Grounding DINO SwinB 0.88 0.84 0.47 0.82 0.76 0.52 0.84 0.80 0.54 0.91 0.89 0.32 0.95 0.94 0.16
SwinT 0.85 0.80 0.53 0.79 0.73 0.60 0.83 0.79 0.59 0.92 0.90 0.28 0.96 0.95 0.16
Bravo-Synobj
CLIP RN50 0.87 0.89 0.65 0.87 0.89 0.65 0.87 0.90 0.71 0.89 0.92 0.66 0.92 0.94 0.57
RN101 0.89 0.92 0.54 0.89 0.92 0.54 0.89 0.92 0.60 0.91 0.93 0.55 0.93 0.95 0.49
RN50x4 0.90 0.92 0.52 0.90 0.92 0.52 0.91 0.93 0.58 0.93 0.95 0.48 0.94 0.95 0.42
RN50x16 0.92 0.94 0.48 0.92 0.94 0.49 0.92 0.94 0.50 0.93 0.94 0.49 0.94 0.96 0.49
RN50x64 0.93 0.94 0.54 0.93 0.94 0.54 0.93 0.95 0.54 0.94 0.95 0.53 0.94 0.96 0.43
ViT-B/32 0.88 0.90 0.57 0.88 0.90 0.56 0.88 0.90 0.61 0.91 0.93 0.52 0.92 0.94 0.54
ViT-B/16 0.91 0.93 0.50 0.91 0.93 0.50 0.91 0.93 0.54 0.92 0.94 0.48 0.93 0.95 0.49
ViT-L/14 0.92 0.94 0.48 0.92 0.94 0.48 0.92 0.94 0.53 0.93 0.95 0.46 0.94 0.96 0.50
DINO ViT-S/16 0.79 0.80 0.74 0.79 0.80 0.74 0.81 0.83 0.74 0.85 0.87 0.66 0.88 0.91 0.63
ViT-S/8 0.83 0.85 0.69 0.83 0.85 0.69 0.84 0.87 0.76 0.87 0.90 0.61 0.89 0.91 0.64
ViT-B/16 0.81 0.82 0.69 0.81 0.82 0.69 0.83 0.85 0.75 0.86 0.88 0.60 0.84 0.87 0.80
ViT-B/8 0.82 0.84 0.64 0.82 0.84 0.64 0.83 0.86 0.70 0.86 0.89 0.59 0.88 0.91 0.65
RN50 0.55 0.54 0.92 0.55 0.54 0.92 0.50 0.50 0.92 0.55 0.54 0.94 0.58 0.57 0.89
DINOv2 ViT-S/14 0.75 0.77 0.77 0.75 0.77 0.77 0.75 0.78 0.82 0.79 0.82 0.77 0.85 0.88 0.74
ViT-B/14 0.79 0.80 0.71 0.79 0.80 0.71 0.80 0.82 0.73 0.84 0.86 0.64 0.87 0.89 0.67
ViT-L/14 0.78 0.79 0.77 0.78 0.79 0.77 0.78 0.81 0.80 0.82 0.85 0.74 0.86 0.88 0.63
ViT-G/14 0.79 0.80 0.73 0.79 0.80 0.73 0.81 0.83 0.69 0.82 0.85 0.71 0.86 0.89 0.69
Grounding DINO SwinB 0.60 0.56 0.83 0.61 0.56 0.82 0.58 0.54 0.85 0.76 0.68 0.58 0.99 0.99 0.02
SwinT 0.57 0.54 0.89 0.58 0.54 0.90 0.57 0.54 0.91 0.64 0.58 0.78 0.91 0.84 0.20
Table 10: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Lost and Found [59] with different intensity values and Bravo-Synobj [50] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
Table 10 presents an extensive evaluation complementing our semantic shift analysis from the main paper. encompassing 19 backbone architectures across four VFMs. This comprehensive comparison spans various architectures including ResNet variants (RN50-RN50x64). Vision Transformers (ViT-B/32 to ViT-L/14). and Swin Transformers.
In addition to the findings presented in the main paper. our comprehensive backbone analysis reveals several key insights discussed in the following in more detail:
Detection Method Sensitivity:
While the main paper highlights the superiority of flow-based methods with Grounding DINO. the full results demonstrate that different VFMs exhibit distinct preferences for detection methods. For instance. CLIP models show remarkable consistency across all four detection methods (APS. OC-SVM. GMM. NF). with RN50x64 maintaining FPR values between 0.37-0.24 on Lost and Found. In contrast. Grounding DINO shows high method sensitivity. with performance varying substantially between GMM (FPR: 0.32) and NF (FPR: 0.16).
Cross-Dataset Generalization:
The extended analysis reveals interesting patterns in cross-dataset performance. Models that excel on real-world data don’t necessarily maintain their advantage on synthetic data. This is particularly evident in CLIP’s RN50x64 performance, which dominates on Lost and Found (AUROC: 0.94) but shows relatively lower performance on Bravo-Synobj (AUROC: 0.93). Conversely. some architectures like CLIP’s ViT-L/14 demonstrate more balanced performance across both datasets. suggesting better generalization capabilities.
Resolution Impact:
The full backbone comparison enables analysis of patch size effects in transformer-based architectures. In DINO models, smaller patch sizes (ViT-B/8 vs ViT-B/16) generally lead to improved performance. Particularly evident in GMM results (AUROC: 0.87 vs 0.84 on Lost and Found). This suggests that finer-grained feature extraction benefits semantic shift detection. though at the cost of computational complexity.
D.2 Covariate Shift Experiments
Model Backbone Foggy Cityscapes (i=0.05)
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.66 0.62 0.80 0.66 0.62 0.80 0.66 0.63 0.84 0.65 0.61 0.81 0.78 0.74 0.62
RN101 0.68 0.65 0.80 0.68 0.65 0.80 0.69 0.68 0.79 0.76 0.73 0.65 0.80 0.75 0.57
RN50x4 0.70 0.66 0.78 0.70 0.66 0.78 0.71 0.68 0.78 0.74 0.70 0.71 0.81 0.77 0.59
RN50x16 0.66 0.62 0.79 0.66 0.62 0.79 0.67 0.63 0.82 0.69 0.65 0.73 0.73 0.67 0.69
RN50x64 0.66 0.62 0.78 0.66 0.62 0.77 0.68 0.64 0.79 0.71 0.66 0.76 0.77 0.73 0.67
ViT-B/32 0.75 0.72 0.64 0.75 0.71 0.65 0.77 0.75 0.62 0.79 0.75 0.58 0.86 0.82 0.49
ViT-B/16 0.73 0.69 0.71 0.73 0.69 0.70 0.74 0.71 0.78 0.78 0.75 0.57 0.86 0.83 0.48
ViT-L/14 0.80 0.76 0.58 0.80 0.76 0.58 0.83 0.80 0.60 0.84 0.82 0.49 0.86 0.82 0.46
DINO ViT-S/16 0.66 0.60 0.68 0.66 0.60 0.68 0.72 0.66 0.68 0.69 0.63 0.65 0.76 0.71 0.60
ViT-S/8 0.61 0.58 0.80 0.61 0.58 0.80 0.65 0.60 0.77 0.65 0.61 0.78 0.73 0.66 0.62
ViT-B/16 0.65 0.60 0.73 0.65 0.60 0.73 0.71 0.65 0.69 0.67 0.63 0.72 0.75 0.72 0.74
ViT-B/8 0.65 0.58 0.72 0.65 0.58 0.72 0.71 0.65 0.68 0.66 0.62 0.72 0.80 0.74 0.52
RN50 0.59 0.56 0.88 0.59 0.56 0.88 0.58 0.57 0.90 0.60 0.57 0.85 0.57 0.56 0.91
DINOv2 ViT-S/14 0.55 0.53 0.90 0.55 0.53 0.90 0.56 0.53 0.90 0.57 0.54 0.86 0.63 0.59 0.83
ViT-B/14 0.58 0.55 0.88 0.58 0.55 0.88 0.60 0.57 0.92 0.59 0.56 0.85 0.66 0.61 0.77
ViT-L/14 0.55 0.53 0.91 0.55 0.53 0.91 0.56 0.54 0.89 0.56 0.53 0.92 0.59 0.55 0.88
ViT-G/14 0.56 0.53 0.89 0.56 0.53 0.89 0.61 0.57 0.84 0.57 0.54 0.89 0.64 0.60 0.83
Grounding DINO SwinB 0.58 0.54 0.85 0.57 0.54 0.84 0.60 0.57 0.86 0.66 0.61 0.79 0.86 0.81 0.36
SwinT 0.57 0.54 0.87 0.57 0.54 0.86 0.59 0.56 0.86 0.69 0.62 0.72 0.84 0.77 0.42
Foggy Cityscapes (i=0.2)
CLIP RN50 0.85 0.82 0.48 0.85 0.82 0.47 0.86 0.84 0.53 0.85 0.81 0.47 0.94 0.91 0.23
RN101 0.87 0.84 0.43 0.87 0.84 0.44 0.88 0.88 0.47 0.94 0.93 0.26 0.94 0.92 0.20
RN50x4 0.90 0.87 0.36 0.90 0.87 0.36 0.91 0.89 0.37 0.94 0.93 0.26 0.96 0.95 0.18
RN50x16 0.88 0.86 0.40 0.88 0.86 0.40 0.89 0.87 0.42 0.90 0.88 0.31 0.92 0.88 0.29
RN50x64 0.90 0.86 0.30 0.90 0.86 0.31 0.91 0.89 0.32 0.92 0.91 0.26 0.94 0.92 0.21
ViT-B/32 0.92 0.91 0.30 0.92 0.90 0.30 0.93 0.93 0.27 0.96 0.95 0.16 0.98 0.98 0.08
ViT-B/16 0.92 0.91 0.29 0.92 0.91 0.29 0.93 0.92 0.29 0.96 0.96 0.18 0.98 0.98 0.10
ViT-L/14 0.96 0.95 0.14 0.96 0.95 0.15 0.97 0.97 0.11 0.98 0.97 0.07 0.97 0.96 0.10
DINO ViT-S/16 0.87 0.81 0.31 0.87 0.81 0.31 0.93 0.90 0.20 0.92 0.89 0.24 0.95 0.94 0.19
ViT-S/8 0.77 0.70 0.50 0.77 0.70 0.50 0.83 0.78 0.45 0.83 0.78 0.44 0.89 0.84 0.30
ViT-B/16 0.85 0.78 0.34 0.85 0.78 0.34 0.91 0.88 0.22 0.89 0.85 0.31 0.89 0.88 0.34
ViT-B/8 0.84 0.77 0.34 0.84 0.77 0.34 0.91 0.87 0.24 0.88 0.85 0.35 0.97 0.96 0.09
RN50 0.73 0.68 0.70 0.73 0.68 0.70 0.70 0.67 0.79 0.73 0.67 0.73 0.65 0.63 0.84
DINOv2 ViT-S/14 0.66 0.60 0.72 0.66 0.60 0.72 0.69 0.62 0.73 0.70 0.63 0.66 0.82 0.76 0.52
ViT-B/14 0.69 0.63 0.73 0.69 0.63 0.73 0.71 0.66 0.77 0.74 0.68 0.66 0.84 0.77 0.48
ViT-L/14 0.62 0.57 0.85 0.62 0.57 0.85 0.62 0.58 0.86 0.63 0.59 0.84 0.69 0.62 0.70
ViT-G/14 0.64 0.58 0.76 0.64 0.58 0.76 0.71 0.64 0.72 0.68 0.62 0.73 0.80 0.74 0.54
Grounding DINO SwinB 0.84 0.79 0.38 0.84 0.79 0.38 0.89 0.86 0.32 0.95 0.94 0.17 1.00 1.00 0.00
SwinT 0.87 0.84 0.34 0.87 0.84 0.34 0.90 0.89 0.30 0.96 0.94 0.10 1.00 1.00 0.02
Bravo-Synflare
CLIP RN50 0.80 0.81 0.67 0.80 0.81 0.68 0.79 0.80 0.77 0.85 0.86 0.54 0.89 0.90 0.42
RN101 0.84 0.85 0.62 0.84 0.85 0.64 0.84 0.86 0.69 0.89 0.91 0.57 0.92 0.93 0.43
RN50x4 0.78 0.79 0.71 0.78 0.79 0.72 0.77 0.79 0.80 0.85 0.86 0.58 0.88 0.87 0.51
RN50x16 0.80 0.81 0.70 0.80 0.81 0.70 0.79 0.80 0.73 0.83 0.84 0.62 0.87 0.88 0.55
RN50x64 0.79 0.79 0.65 0.79 0.79 0.63 0.81 0.81 0.63 0.82 0.83 0.62 0.85 0.82 0.47
ViT-B/32 0.84 0.84 0.60 0.84 0.84 0.61 0.83 0.84 0.62 0.91 0.91 0.47 0.93 0.94 0.37
ViT-B/16 0.84 0.85 0.52 0.84 0.84 0.53 0.83 0.84 0.60 0.89 0.90 0.43 0.94 0.95 0.35
ViT-L/14 0.87 0.86 0.49 0.87 0.86 0.50 0.88 0.89 0.48 0.90 0.91 0.45 0.94 0.93 0.31
DINO ViT-S/16 0.86 0.85 0.47 0.86 0.85 0.47 0.90 0.90 0.43 0.91 0.91 0.42 0.95 0.96 0.28
ViT-S/8 0.82 0.81 0.55 0.82 0.81 0.55 0.86 0.85 0.51 0.87 0.88 0.54 0.93 0.93 0.35
ViT-B/16 0.87 0.87 0.41 0.87 0.87 0.41 0.91 0.91 0.33 0.91 0.92 0.37 0.87 0.88 0.57
ViT-B/8 0.87 0.87 0.45 0.87 0.87 0.45 0.91 0.92 0.35 0.90 0.91 0.51 0.95 0.96 0.33
RN50 0.64 0.60 0.84 0.64 0.60 0.84 0.54 0.53 0.90 0.69 0.67 0.82 0.64 0.62 0.87
DINOv2 ViT-S/14 0.70 0.67 0.80 0.70 0.67 0.80 0.69 0.67 0.88 0.76 0.74 0.78 0.79 0.76 0.72
ViT-B/14 0.64 0.62 0.86 0.64 0.62 0.86 0.64 0.62 0.87 0.70 0.68 0.85 0.76 0.75 0.78
ViT-L/14 0.60 0.58 0.88 0.60 0.58 0.88 0.60 0.58 0.89 0.61 0.61 0.88 0.67 0.65 0.80
ViT-G/14 0.60 0.57 0.83 0.60 0.57 0.83 0.64 0.61 0.83 0.63 0.62 0.88 0.72 0.72 0.80
Grounding DINO SwinB 0.86 0.85 0.51 0.86 0.85 0.51 0.86 0.86 0.49 0.95 0.95 0.26 0.99 0.99 0.05
SwinT 0.85 0.84 0.49 0.85 0.84 0.49 0.84 0.84 0.55 0.94 0.94 0.25 0.99 0.99 0.07
Table 11: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Foggy Cityscapes [59] with different intensity values and Bravo-Synflare [50] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
Tabs. 11, 12, 13 and 14 present comprehensive covariate shift experiments that extend the analysis introduced in the main paper. Specifically, Tabs. 11 and 12 demonstrate experimental results where the monitoring set $\mathcal{X}_{\mathrm{monitor}}$ is derived from the Cityscapes training set, while Tabs. 13 and 14 utilize the ACDC reference training set as $\mathcal{X}_{\mathrm{monitor}}$.
Extending beyond the main paper’s findings, our comprehensive backbone analysis reveals several key insights:
Performance Variation in Environmental Conditions:
Analysis of Grounding DINO reveals a notable performance discrepancy between synthetic and natural environmental shifts. While achieving near-perfect detection on synthetic perturbations (FPR95: 0.00-0.02 on Foggy Cityscapes i=0.2), performance degrades on natural conditions, particularly in ACDC Rain (FPR95: 0.34-0.35). This degradation is especially interesting as ACDC Rain represents post-rain conditions (e.g., wet road surfaces, residual moisture) rather than active precipitation. .
Architecture-Specific Scaling Patterns:
Our analysis reveals non-linear scaling behaviors across architectural families. Within the CLIP family, model size does not correlate monotonically with performance. While ViT-L/14 demonstrates superior performance on synthetic perturbations in Foggy Cityscapes, this advantage diminishes or reverses in natural conditions such as ACDC Rain, where RN50x64 achieves markedly better performance (FPR95: 0.17 vs 0.41). These findings suggest that architectural design choices may be more crucial than model capacity for robust covariate shift detection.
Model Backbone Bravo-Synrain
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.76 0.72 0.59 0.76 0.72 0.60 0.75 0.71 0.64 0.85 0.82 0.43 0.95 0.92 0.16
RN101 0.83 0.81 0.50 0.83 0.81 0.50 0.84 0.82 0.51 0.92 0.91 0.25 0.94 0.93 0.17
RN50x4 0.85 0.82 0.48 0.85 0.82 0.48 0.85 0.82 0.48 0.93 0.92 0.24 0.95 0.93 0.14
RN50x16 0.94 0.93 0.25 0.94 0.93 0.25 0.94 0.94 0.22 0.96 0.95 0.17 0.97 0.96 0.10
RN50x64 0.99 0.99 0.03 0.99 0.99 0.03 1.00 1.00 0.01 1.00 1.00 0.01 0.99 0.99 0.04
ViT-B/32 0.82 0.79 0.53 0.82 0.79 0.54 0.82 0.80 0.50 0.91 0.89 0.31 0.96 0.95 0.13
ViT-B/16 0.90 0.87 0.34 0.90 0.87 0.34 0.91 0.89 0.31 0.94 0.93 0.20 0.97 0.95 0.12
ViT-L/14 0.89 0.85 0.31 0.89 0.85 0.32 0.92 0.90 0.27 0.94 0.91 0.22 0.97 0.94 0.09
DINO ViT-S/16 0.87 0.83 0.33 0.87 0.83 0.33 0.93 0.91 0.20 0.94 0.93 0.21 1.00 1.00 0.02
ViT-S/8 0.89 0.84 0.30 0.89 0.84 0.30 0.95 0.93 0.16 0.96 0.95 0.13 0.99 0.99 0.02
ViT-B/16 0.90 0.87 0.25 0.90 0.87 0.25 0.96 0.95 0.11 0.95 0.94 0.17 0.99 0.99 0.03
ViT-B/8 0.95 0.94 0.15 0.95 0.94 0.15 0.99 0.99 0.06 0.98 0.98 0.10 1.00 1.00 0.00
RN50 0.49 0.47 0.92 0.49 0.47 0.92 0.40 0.42 0.97 0.56 0.52 0.87 0.65 0.64 0.85
DINOv2 ViT-S/14 0.89 0.83 0.33 0.89 0.83 0.33 0.89 0.84 0.35 0.96 0.92 0.16 0.91 0.84 0.20
ViT-B/14 0.80 0.75 0.58 0.80 0.75 0.58 0.81 0.76 0.61 0.88 0.83 0.45 0.94 0.90 0.18
ViT-L/14 0.73 0.67 0.64 0.73 0.67 0.64 0.74 0.68 0.67 0.80 0.75 0.56 0.88 0.83 0.37
ViT-G/14 0.71 0.65 0.66 0.71 0.65 0.66 0.77 0.71 0.57 0.79 0.74 0.61 0.90 0.87 0.35
Grounding DINO SwinB 1.00 1.00 0.02 1.00 1.00 0.02 1.00 0.99 0.02 1.00 1.00 0.00 1.00 1.00 0.00
SwinT 0.99 0.99 0.02 0.99 0.99 0.02 0.99 0.99 0.03 1.00 1.00 0.00 1.00 1.00 0.00
Table 12: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on Bravo-Synrain [50] dataset. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
Model Backbone ACDC Fog
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.95 0.94 0.22 0.95 0.94 0.22 0.94 0.93 0.22 0.98 0.98 0.07 0.97 0.94 0.09
RN101 0.94 0.94 0.25 0.94 0.94 0.25 0.94 0.93 0.28 0.98 0.97 0.12 0.98 0.96 0.09
RN50x4 0.97 0.97 0.13 0.97 0.97 0.13 0.97 0.96 0.16 0.99 0.98 0.07 0.98 0.98 0.06
RN50x16 0.97 0.97 0.11 0.97 0.97 0.11 0.97 0.97 0.12 0.98 0.98 0.08 0.98 0.96 0.08
RN50x64 0.98 0.98 0.07 0.98 0.98 0.07 0.99 0.99 0.06 0.98 0.98 0.06 0.99 0.98 0.04
ViT-B/32 0.97 0.97 0.12 0.97 0.97 0.12 0.97 0.97 0.11 0.99 0.99 0.03 0.99 0.99 0.04
ViT-B/16 0.95 0.95 0.20 0.95 0.95 0.20 0.95 0.95 0.18 0.98 0.98 0.07 0.98 0.98 0.07
ViT-L/14 0.98 0.97 0.09 0.98 0.97 0.09 0.98 0.97 0.09 1.00 0.99 0.02 0.99 0.98 0.03
DINO ViT-S/16 0.80 0.77 0.57 0.80 0.77 0.57 0.79 0.76 0.60 0.90 0.88 0.34 0.87 0.84 0.44
ViT-S/8 0.68 0.63 0.71 0.68 0.63 0.71 0.70 0.65 0.70 0.77 0.70 0.56 0.81 0.77 0.52
ViT-B/16 0.74 0.70 0.66 0.74 0.70 0.66 0.74 0.70 0.65 0.86 0.82 0.43 0.85 0.83 0.50
ViT-B/8 0.81 0.77 0.48 0.81 0.77 0.48 0.83 0.80 0.46 0.90 0.88 0.35 0.91 0.90 0.35
RN50 0.92 0.91 0.42 0.92 0.91 0.42 0.92 0.92 0.38 0.90 0.90 0.38 0.64 0.65 0.89
DINOv2 ViT-S/14 0.82 0.79 0.54 0.82 0.79 0.54 0.79 0.75 0.58 0.91 0.87 0.23 0.90 0.83 0.24
ViT-B/14 0.74 0.71 0.67 0.74 0.71 0.67 0.73 0.68 0.69 0.84 0.78 0.46 0.85 0.79 0.44
ViT-L/14 0.61 0.58 0.84 0.61 0.58 0.84 0.60 0.56 0.79 0.72 0.66 0.69 0.75 0.67 0.59
ViT-G/14 0.62 0.60 0.82 0.62 0.60 0.82 0.67 0.63 0.71 0.77 0.73 0.63 0.83 0.77 0.46
Grounding DINO SwinB 0.76 0.73 0.57 0.75 0.72 0.58 0.82 0.81 0.50 0.96 0.95 0.16 0.98 0.97 0.08
SwinT 0.80 0.79 0.54 0.80 0.78 0.55 0.83 0.83 0.52 0.98 0.97 0.08 0.96 0.95 0.13
ACDC Night
CLIP RN50 0.94 0.91 0.17 0.94 0.91 0.17 0.92 0.89 0.23 0.93 0.90 0.27 0.92 0.84 0.19
RN101 0.96 0.94 0.15 0.96 0.94 0.16 0.92 0.89 0.25 0.87 0.83 0.39 0.91 0.84 0.22
RN50x4 0.98 0.98 0.07 0.98 0.98 0.07 0.94 0.93 0.19 0.87 0.83 0.43 0.83 0.75 0.34
RN50x16 0.97 0.96 0.09 0.97 0.96 0.10 0.95 0.93 0.16 0.85 0.82 0.46 0.88 0.78 0.27
RN50x64 0.97 0.96 0.09 0.97 0.96 0.09 0.96 0.95 0.12 0.96 0.94 0.13 0.90 0.84 0.23
ViT-B/32 0.90 0.88 0.32 0.90 0.88 0.32 0.87 0.84 0.37 0.91 0.88 0.27 0.90 0.84 0.23
ViT-B/16 0.91 0.89 0.25 0.91 0.89 0.26 0.89 0.85 0.31 0.88 0.85 0.34 0.93 0.88 0.16
ViT-L/14 0.91 0.88 0.27 0.91 0.88 0.27 0.88 0.85 0.38 0.86 0.82 0.35 0.92 0.86 0.20
DINO ViT-S/16 0.90 0.85 0.23 0.90 0.85 0.23 0.89 0.83 0.28 0.82 0.76 0.45 0.87 0.82 0.36
ViT-S/8 0.85 0.77 0.31 0.85 0.77 0.31 0.85 0.77 0.32 0.76 0.67 0.48 0.82 0.78 0.50
ViT-B/16 0.88 0.82 0.26 0.88 0.82 0.26 0.88 0.82 0.28 0.87 0.82 0.33 0.94 0.92 0.23
ViT-B/8 0.93 0.88 0.18 0.93 0.88 0.18 0.92 0.87 0.20 0.78 0.71 0.48 0.91 0.88 0.27
RN50 0.90 0.87 0.34 0.90 0.87 0.34 0.84 0.82 0.57 0.92 0.90 0.26 0.71 0.68 0.76
DINOv2 ViT-S/14 0.89 0.83 0.25 0.89 0.83 0.25 0.85 0.78 0.35 0.77 0.71 0.63 0.79 0.70 0.53
ViT-B/14 0.86 0.79 0.35 0.86 0.79 0.35 0.85 0.78 0.39 0.85 0.79 0.40 0.82 0.75 0.46
ViT-L/14 0.84 0.78 0.42 0.84 0.78 0.42 0.85 0.79 0.39 0.84 0.78 0.48 0.85 0.79 0.39
ViT-G/14 0.87 0.81 0.35 0.87 0.81 0.35 0.88 0.83 0.31 0.89 0.86 0.31 0.89 0.86 0.31
Grounding DINO SwinB 0.96 0.95 0.13 0.96 0.95 0.13 0.97 0.96 0.13 0.98 0.97 0.10 0.99 0.99 0.02
SwinT 0.96 0.95 0.13 0.96 0.95 0.13 0.96 0.95 0.14 0.99 0.99 0.04 0.98 0.98 0.08
Table 13: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on ACDC Fog and ACDC Night [69] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
Model Backbone ACDC Rain
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.89 0.86 0.36 0.89 0.86 0.37 0.90 0.87 0.32 0.93 0.92 0.24 0.94 0.90 0.21
RN101 0.87 0.84 0.42 0.87 0.84 0.42 0.87 0.84 0.39 0.91 0.89 0.30 0.91 0.87 0.25
RN50x4 0.89 0.86 0.36 0.89 0.86 0.36 0.90 0.88 0.30 0.93 0.92 0.22 0.93 0.89 0.21
RN50x16 0.89 0.86 0.31 0.89 0.86 0.31 0.90 0.87 0.28 0.91 0.89 0.28 0.93 0.89 0.19
RN50x64 0.95 0.94 0.17 0.95 0.94 0.17 0.96 0.96 0.14 0.92 0.90 0.26 0.92 0.87 0.22
ViT-B/32 0.85 0.83 0.41 0.85 0.83 0.41 0.87 0.84 0.39 0.92 0.90 0.26 0.95 0.92 0.18
ViT-B/16 0.84 0.79 0.44 0.84 0.79 0.44 0.86 0.81 0.38 0.91 0.89 0.27 0.90 0.85 0.25
ViT-L/14 0.84 0.78 0.41 0.84 0.78 0.42 0.87 0.81 0.35 0.86 0.83 0.42 0.92 0.87 0.23
DINO ViT-S/16 0.78 0.73 0.64 0.78 0.73 0.64 0.80 0.76 0.60 0.82 0.80 0.61 0.84 0.84 0.52
ViT-S/8 0.80 0.77 0.62 0.80 0.77 0.62 0.82 0.80 0.54 0.83 0.83 0.55 0.86 0.87 0.56
ViT-B/16 0.81 0.77 0.56 0.81 0.77 0.56 0.84 0.80 0.53 0.85 0.84 0.53 0.81 0.83 0.68
ViT-B/8 0.82 0.80 0.51 0.82 0.80 0.51 0.85 0.82 0.47 0.84 0.84 0.52 0.84 0.86 0.58
RN50 0.66 0.62 0.78 0.66 0.62 0.78 0.59 0.57 0.88 0.70 0.67 0.81 0.55 0.54 0.91
DINOv2 ViT-S/14 0.75 0.71 0.69 0.75 0.71 0.69 0.78 0.74 0.70 0.79 0.78 0.66 0.84 0.82 0.62
ViT-B/14 0.70 0.67 0.74 0.70 0.67 0.74 0.71 0.68 0.71 0.71 0.70 0.78 0.73 0.69 0.75
ViT-L/14 0.63 0.61 0.80 0.63 0.61 0.80 0.67 0.65 0.79 0.67 0.66 0.83 0.75 0.72 0.67
ViT-G/14 0.64 0.61 0.77 0.64 0.61 0.77 0.71 0.67 0.69 0.72 0.69 0.71 0.78 0.73 0.58
Grounding DINO SwinB 0.75 0.70 0.63 0.76 0.71 0.62 0.72 0.66 0.67 0.77 0.73 0.60 0.89 0.85 0.35
SwinT 0.71 0.66 0.71 0.71 0.66 0.70 0.67 0.62 0.68 0.81 0.76 0.52 0.89 0.86 0.34
ACDC Snow
CLIP RN50 0.98 0.97 0.08 0.98 0.97 0.08 0.98 0.97 0.08 0.99 0.99 0.05 0.98 0.96 0.08
RN101 0.96 0.94 0.14 0.96 0.94 0.14 0.96 0.95 0.12 0.98 0.97 0.09 0.98 0.97 0.06
RN50x4 0.98 0.97 0.08 0.98 0.97 0.08 0.98 0.98 0.06 0.99 0.99 0.07 0.99 0.99 0.04
RN50x16 0.98 0.97 0.07 0.98 0.97 0.07 0.98 0.97 0.06 0.99 0.99 0.04 0.98 0.96 0.05
RN50x64 0.99 0.98 0.05 0.99 0.98 0.05 0.99 0.99 0.04 0.99 0.98 0.05 0.99 0.98 0.03
ViT-B/32 0.96 0.96 0.12 0.96 0.96 0.12 0.97 0.96 0.11 0.99 0.99 0.05 0.99 0.99 0.03
ViT-B/16 0.95 0.94 0.16 0.95 0.94 0.16 0.95 0.94 0.15 0.99 0.99 0.05 0.99 0.98 0.05
ViT-L/14 0.93 0.90 0.20 0.93 0.90 0.20 0.94 0.92 0.15 0.97 0.97 0.11 0.98 0.97 0.05
DINO ViT-S/16 0.81 0.76 0.46 0.81 0.76 0.46 0.83 0.79 0.45 0.87 0.84 0.43 0.86 0.83 0.51
ViT-S/8 0.76 0.70 0.58 0.76 0.70 0.58 0.79 0.73 0.52 0.81 0.76 0.48 0.89 0.87 0.41
ViT-B/16 0.82 0.76 0.48 0.82 0.76 0.48 0.84 0.79 0.42 0.86 0.83 0.43 0.88 0.86 0.48
ViT-B/8 0.83 0.78 0.40 0.83 0.78 0.40 0.87 0.82 0.33 0.89 0.86 0.38 0.92 0.92 0.34
RN50 0.80 0.76 0.59 0.80 0.76 0.59 0.74 0.72 0.72 0.82 0.78 0.59 0.61 0.60 0.90
DINOv2 ViT-S/14 0.91 0.86 0.22 0.91 0.86 0.22 0.94 0.90 0.18 0.94 0.93 0.28 0.96 0.93 0.15
ViT-B/14 0.81 0.76 0.51 0.81 0.76 0.51 0.85 0.80 0.43 0.87 0.85 0.49 0.94 0.91 0.23
ViT-L/14 0.74 0.67 0.63 0.74 0.67 0.63 0.78 0.72 0.55 0.79 0.74 0.62 0.88 0.83 0.38
ViT-G/14 0.71 0.65 0.66 0.71 0.65 0.66 0.78 0.72 0.48 0.82 0.77 0.49 0.91 0.86 0.28
Grounding DINO SwinB 0.79 0.75 0.59 0.79 0.75 0.58 0.80 0.75 0.52 0.93 0.92 0.24 0.99 0.98 0.04
SwinT 0.76 0.71 0.59 0.76 0.71 0.60 0.75 0.70 0.59 0.94 0.92 0.24 0.98 0.97 0.07
Table 14: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on ACDC Rain and ACDC Snow [69] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
Appendix E Limitations and Future Directions
Utilizing text-encoders:
Our current work focuses exclusively on leveraging the image encoding capabilities of VFMs. However, many VFMs, such as Grounding-DINO [45] and CLIP [61] feature a shared latent space enabling cross-modal alignment between image and text representations. Our approach currently utilizes only image-level representations to detect distribution shifts. As a result, our approach primarily addresses distribution shifts at the image level, leaving the potential of text-based representations unexplored.
Incorporating text representations could provide richer characterization of different types of distribution shifts, where the text encoder could generate human-interpretable descriptions of both semantic and covariate shifts, offering valuable insights for subsequent mitigation strategies and system interpretability.
Combination with Pixel-wise OOD Segmentation:
Recent advances by Ren et al. [64] demonstrate the feasibility of real-time deployment through Grounding-DINO Edge, suggesting that dual-encoder VFMs can be effectively integrated into the operational domain of AD systems. This integration of real-time cross-modal analysis presents a promising direction for future research in robust perception systems, potentially enabling more comprehensive and interpretable OOD detection.
Towards Treatment of OOD Samples:
This work solely concentrated on the hard problem to reliably detect OOD input images. However, the benefit of the monitor only becomes practically relevant, if it is combined with subsequent mitigation techniques that handle OOD cases. Our experiments assumed perfect mitigation techniques (we simply filtered OOD samples out), equivalent to a 100% reliable handover to a 100% reliable backup driver. Apart from handovers, alternative backup routines could be, e.g., to apply a safe(r) state (e.g., get to a hold), or query a potentially more expensive expert model for which better results are expected. For future work it will be highly interesting to see whether the VFM OOD detectors provide valuable information about the OOD samples that can be leveraged for the OOD case treatment.